### v1

Abstract
The welfare of a dynasty that consists of identical Bernoullian utilities subject to Cobb-Douglas production with constant labor is maximized. We blend two discounting schemes to avoid the objective function from exploding to infinity: one is the compound discounting of future utilities and the other is the binary discounting which undiscounts utilities before the planning horizon but fully discounts those that lie beyond. After presenting a general solution to the optimal schedule of consumption, we study the planning solution analytically and also numerically, to discover that the optimal planning horizon is not necessarily an infinity, even under the case of zero compound discounting. Furthermore, we study the inequality of consumptions among generations with respect to the compound discount rates and planning horizons.
[…]
X00 Y00
Keywords:
Dynamic Programming Sustainability Optimal Population Repugnant Conclusion Social Discounting Intergenerational Equity
journal:
1 Introduction
This study is concerned with population axiology from a dynamical perspective. Greaves (2017) delineates population axiology as the study of the conditions under which one state of affairs is better than another, when the states of affairs in question may differ over the numbers and the identities of the persons. There are several approaches to population axiology; one most commonly employed in economics is the total view (utilitarianism) that evaluates populations according to the total amount of well-being that they contain. Among the variety of principles of aggregating potential well-beings, a critical-level utilitarian (e.g., Blackorby et al, 2005) particularly seeks for a state $x$ to maximize the population value $\mathcal{\mathcal{V}}=\sum_{i\in n(x)}\left({\Upsilon}_{i}(x)-\kappa\right)$, where $n(x)$ is the set of individuals, ${\Upsilon}_{i}(x)$ is the lifetime utility of an individual $i$, and $\kappa\geq 0$ is the critical level.
[…]
Under classical utilitarianism ($\kappa=0$), $s=\bar{c}>0$ is called the well-being subsistence (Dasgupta, 2019) or consumption for neutrality, since ${\Upsilon}(s)=0$. For any consumption $c_{i}<s$, utility level of an individual $i$ is negative, meaning that $i$’s life is not worth living. Hence, it must be that $c_{i}>s$ so the lifetime utility of all individuals are positive. Note however that under critical-level utilitarianism, an individual $i$’s life contributes positively to the value of a population if $\bar{c}<c_{i}$, while negatively if $s<c_{i}<\bar{c}$, even the person leads a life worth living. For analytical purposes, we employ ${\Upsilon}(c_{i})=\log c_{i}+\kappa$ with a $\kappa$ that the well-being subsistence $s=e^{-\kappa}$ is sufficiently close to zero. The population value of critical-level utilitarianism will then be $\mathcal{\mathcal{V}}=\sum_{i=1}^{n}\log c_{i}$ where $c_{i}\in[e^{-\kappa},\infty)$ and $\bar{c}=1$, which we hereafter seek to maximize, with respect to $c_{i}$ and $n$.
As we recall our primary question, let us introduce descrete time in population ethics while assuming, for sake of simplicity, that periodical number of people is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ they live in. [^2] We may therefore call $t$ a generation. The population ethics problem then becomes a generation ethics problem. The critical-level utilitarian population value function now becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$. We note that this type of value function (based on Bernoullian utility) is widely employed in the previous Ramsey-Cass-Koopmans type of dynastic optimization models (e.g., Nordhaus, 1992; Stern, 2007), except that the future generations are discounted according to the periodical deviation from the present, and typically, the number of potential generations $n$ is left unquestioned (and hence, set to infinity). In contrast, potential generations will be accounted equally whilst the number of generations will be subject to maximization, in this study.
The remainder of the paper proceeds as follows. In the following section, we first specify the model of a production economy based on Cobb-Douglas technology from a dynamical perspective. We then specify our primary question in the form of a finite horizon dynamical programming problem where the objective is to maximize the sum of utilities given the abovementioned model of production, and derive a general solution function. In Section 3 we introduce two broad settings of parameters to simplify the general solution function. The first one is the AK setting where the output elasticity of capital is set to unity. The second one is the ZD (zero discounting) setting where the discount factor is set to unity. We also study the peculiarity of the two models in regard to intergenerational equity and optimal initial action of consumption. Section 4 concludes the paper.
[…]
Figure 1: Contribution to the population value (vertical axis) under critical-level utilitarianism. Left: The critical-level consumption $\bar{c}$ contributes none to the population value, i.e., ${\Upsilon}(\bar{c})-\kappa=0$. Consumption level $c^{*}$ maximizes the critical-level utilitarian population value. Right: Well-being subsistence $s$ bears zero lifetime utility (neutrality), i.e., ${\Upsilon}(s)=0$.
2 The model
2.1 The problem of a dynasity
Let us begin by postulating a two-factor Cobb-Douglas production function of constant returns to scale, as follows:
$Y_{t}={A}(K_{t})^{\theta}(L_{t})^{1-\theta}=\frac{{B}}{\theta^{\theta}(1-\theta)^{1-\theta}}(K_{t})^{\theta}(L_{t})^{1-\theta}$ (1)
where $Y_{t}$ denotes total output of the economy, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all effective in period $t$. As for the parameters concerned, $\theta<1$ denotes the output elasticity of capital, ${A}>0$ denotes the level of technology termed productivity, and ${B}={A}\Theta$, where $\Theta=\theta^{\theta}(1-\theta)^{1-\theta}$, denotes the cost function-based productivity, which we otherwise call the dual productivity.333 See Appendix 1 for more details.
Below describes the breakdown of total output into investment, depreciation of capital, and final consumption:
$Y_{t}=K_{t+1}-K_{t}+\delta K_{t}+C_{t}$ (2)
Here, $\delta\leq 1$ denotes depreciation factor of capital, and $C_{t}$ denotes consumption in period $t$. Combining equations (1) and (2), and dividing both sides by $L_{t}=L$ which we assume to be constant over time, leads to the following intertemporal dynamics of capital intensity:
$y_{t}={A}(k_{t})^{\theta}=k_{t+1}-(1-\delta)k_{t}+c_{t}$ (3)
where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in period $t$ and $t+1$, respectively. Also, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per caput output and consumption, respectively, in period $t$.
[…]
2.2 Finite horizon dynamic programming
Our primary problem here is to solve for an optimal consumption path given any planning horizon. The Bellman equation of the problem is of the following:555 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
$\mathcal{\mathcal{V}}_{t}\left[k_{t};\ n\right]=\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{\mathcal{V}}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};\ n\right]\right)$ (4)
The optimal trajectory of the state variable $k_{t}$ is given, by Lemma 2 which we append to Appendix 2 with proofs, as follows:
$k_{t}^{*}[n]=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)^{\theta^{t-i}}=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i-1}}$ (5)
where $S_{\ell}$ is defined as follows:
$S_{\ell}\equiv\sum_{i=0}^{\ell}(\beta\theta)^{i}=1+(\beta\theta)+(\beta\theta)^{2}+\cdots+(\beta\theta)^{\ell}=\frac{1-(\beta\theta)^{\ell+1}}{1-\beta\theta}$ (6)
The following optimal trajectory of consumption is obtained by (5), and (21) which must be true by the proof of Lemma 1.
$c_{t}^{*}[n]=\frac{{A}\left(k_{t}^{*}[n]\right)^{\theta}}{S_{n-1}}=\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{n}-t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i}}$ (7)
With formula (7) the population value function is given as follows:
$\mathcal{V}[n]$ $=\sum_{t=0}^{n}\beta^{t}\log c_{t}^{*}[n]$
$=\log\left(\frac{{A}(k_{0})^{\theta}}{S_{n}}\right)+\sum_{t=1}^{n}\beta^{t}\log\left(\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{n}-t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i}}\right)$ (8)
We hereafter seek to maximize this function with respect to the planning horizon $n$.
Our approach to the analysis of population value (8) will be numerical rather than analytical, because the derivative for $\mathcal{V}[n]$ with respect to $n$ seems uninformative. As it turns out in the following section, the optimal tragectory of contributions $\log c_{t}^{*}[n]$ and the population value $\mathcal{V}[n]$, for any given planning horizon $n$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$. We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon, is evaluable under $\beta\theta<1$. We therefore base our study of zero discounting ($\beta=1$) on this setting (i.e., $\beta\theta<1$, which indicates $\theta<1$, a Cobb-Douglas model). In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$) which we term as AK settings, and Cobb-Douglas production without future discounting ($\theta<1$ and $\beta=1$) which we term as ZD settings.
Table 1 summarizes the parameter settings chosen for numerical examinations. Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log({A}\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log(A\Theta)$. The final case VIII corresponds to the parameter settings that exemplifies $\log({A}\beta)=\log(A\Theta)=0$.
Table 1: Parameters applied in various cases.
case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log({A\Theta})$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.675$
II $1.01$ $0.992$ $1$ $+$ 95 $59.965$
III *1 $0.992$ $1$ 00 73 $55.627$
IV $1.005$ $0.992$ $1$- 58 $50.979$
V $1.05$ $1$ $0.992$ $+$ $\infty$ +$\infty$+ ∞
VI $1.2$ $1$2 00 (281) $41.688$
VII $1.05$ $1$ $0.991$- 117 $\infty$- ∞
VIII $1$ $1$ $1$ 00 00 54 $55.182$$\infty$- ∞
  • Note:
    For all cases I–VIII, $k_{0}=150$.
  • *1
    ${A}=1/\beta\approx 1.008065$.
  • *2
    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=$ 1.
3 Analysis
3.1 AK setting
AK model is one of the simplest yet fundamental models of endogenous growth, which was formally developed by Frankel (1962). Here, we take this production model to study the population value function that discounts future generations to see whether if $n\to\infty$ is an optimal policy. The optimal consumption path (7) for AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
$c_{t}^{*}[n]=\left(\frac{{A}k_{0}}{S_{{n}-t}}\right)\left(\frac{S_{{n}-1}}{S_{n}}{A}\beta\right)\left(\frac{S_{{n}-2}}{S_{{n}-1}}{A}\beta\right)\cdots\left(\frac{S_{{n}-t}}{S_{{n}-t+1}}{A}\beta\right)=\frac{({A}\beta)^{t}{A}k_{0}}{S_{n}}$ (9)
The population value function (8), therefore, becomes:
$\mathcal{V}[n]$ $=\sum_{t=0}^{n}\beta^{t}\log\left(\frac{({A}\beta)^{t}{A}k_{0}}{S_{n}=1+\beta+\cdots+\beta^{n}}\right)$
$=\frac{\beta-((1-\beta){n}+1)\beta^{{n}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}{A}k_{0}\right)$ (10)
For sake of the analysis let us take the derivative with respect to ${n}$.
$\frac{\mathrm{d}\mathcal{V}[n]}{\mathrm{d}n}=\lambda\left(\gamma-{n}\log({A}\beta)+\log\left(1-\beta^{n+1}\right)\right)$ (11)
where $\lambda$ and $\gamma$, that are both independent of $n$, are given as follows:
$\lambda=\frac{\beta^{n}\log\beta^{\beta}}{1-\beta}>0,$ $\gamma=1-\frac{\log({A}\beta)^{1-\beta+\log\beta}}{\log\beta^{1-\beta}}-\log({A}(1-\beta)k_{0})$
In order to study the derivative sign of (11), we consider the following two functions whose equality convey the first order conditions, i.e., $f[n]=g[n]$.
$f[n]=-\gamma+n\log({A}\beta),$ $g[n]=\log\left(1-\beta^{n+1}\right)$ (12)
Before we proceed, let us examine the possible range of the parameter $A$. Described below is the evaluation of the marginal product of capital (MPK) of our production function (1) net of depreciation, which should coincide with the real interest rate $\rho>0$, viz.,
$\frac{\partial Y_{t}}{\partial K_{t}}-\delta={A}\theta\left(\frac{K_{t}}{L_{t}}\right)^{\theta-1}-\delta=\rho>0$ (13)
Then, we know that AK models with complete capital depreciation $\theta=\delta=1$ amounts to ${A}=1+\rho$, and therefore ${A}>1$ is a relevant assumption in this setting.666 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross rate of interest and the rate of generational preference of the population.
Figure 2 (left) depicts functions $f[n]$ and $g[n]$ under different parameters for cases I–IV. Clearly, $f[n]$ is a linear function whose slope is $\log({A}\beta)$, and $g[n]$ is monotonously increasing and approaching zero, i.e., $g[\infty]=0$. Here, we fix the discount factor $\beta$, and differentiate the productivity ${A}$ in four different levels. As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[n]$ will intersect with $g[n]$ at a single point and the population value function $\mathcal{V}[n]$ has a single peak. If ${A}>1/\beta$ so that $f[n]$ has a positive slope, then $f[n]$ and $g[n]$ could either intersect with two points where the population value $\mathcal{V}[n]$ would rise, fall and then rise again, or never intersect so that the population value would be rising indefinitely, with respect to the planning horizon $n$. Figure 2 (right) depicts the population value functions for cases I–IV.
Figure 2: Left: Solid plots show $g[n]$ defined in (12) for cases I–IV with obvious correspondences, i.e., the steepest one corresponds to case I, whereas case IV corresponds to the one with a negative slope. Dashed plot shows $f[n]$ also defined in (12). Right: The plots show population value paths $\mathcal{V}[n]$ for cases I–IV with obvious correspondences.
To visualize the optimal trajectory of the variables in various situations, we specify them here in the form of functionals. By virtue of (9), the optimal trajectory of undiscounted contribution to the population value becomes linear with respect to $t$, as follows:
$\log c_{t}^{*}[n]=\log\left(\frac{1-\beta}{1-\beta^{n+1}}{A}k_{0}\right)+t\log({A}\beta)$ (14)
By virtue of (21), the optimal trajectory of capital intensity becomes as follows:
$k_{t}^{*}[n]=\frac{S_{n-t}}{{A}}c_{t}^{*}[n]=\frac{1-\beta^{n+1-t}}{1-\beta^{n+1}}({A}\beta)^{t}k_{0}$ (15)
Figure 3 displays the population value function $\mathcal{V}[n]$, specified as (10), for cases I, III, and IV, on the top. [^7] These figures are reiterations from Figure 2 (right). Note that, as long as $\beta<1$, the population value function will always converge to some finite value, regardless of the parameter settings besides $\beta$. As we let $n\to\infty$ in (10), we have what follows:
$\mathcal{V}[\infty]=\frac{\log\left({A}(1-\beta)^{1-\beta}\beta^{\beta}\right)}{(1-\beta)^{2}}+\frac{\log(k_{0})}{1-\beta}$ (16)
By comparing the population values at the numerical solution of the first order condition $f[n]=g[n]$ in reference to (12), with $\mathcal{V}[\infty]$ in reference to (16), we know the optimal planning horizon $n^{*}$, which we display in Table 1 for cases I–IV. At the middle row of Figure 3, optimal trajectories of undiscounted contribution to the population value, according to (14), with respect to planning horizons $n=200,400,600$, are displayed for sample I, III, and IV from left to right. Similarly at the bottom row of Figure 3, optimal trajectories of capital intensity, according to (15), with respect to planning horizons $n=200,400,600$, are displayed for cases I, III, and IV from left to right.
Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK model cases I (left), III (center), and IV (right).
3.2 ZD setting
Here, we study zero discounting $\beta=1$ under Cobb-Douglas production $\theta<1$. The condition that $\beta\theta<1$ nonetheless allows us to evaluate the key summation as follows:
$S_{\infty-\tau}=\lim_{n\to\infty}\frac{1-(\beta\theta)^{n-\tau+1}}{1-\beta\theta}=\frac{1}{1-\beta\theta}$
As we apply the above and $n\to\infty$ into (7), we have the following:
$\log c_{t}^{*}[\infty]$ $=\log\left(\frac{{A}}{S_{{\infty}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{\infty}-i}}{S_{{\infty}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
$=\log\left({{A}}{(1-\beta\theta)}\right)+\left(\theta^{t-1}+\cdots+\theta+1\right)\log({A}\beta\theta)^{\theta}+\theta^{t+1}\log k_{0}$
$=\frac{\log\left({A}(1-\beta\theta)^{1-\theta}(\beta\theta)^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\beta\theta)^{\frac{\theta}{1-\theta}}\right)$ (17)
where the optimal trajectory for undiscounted contributions turns out to be geometrically convergent. We can then apply $\beta=1$ to arrive at the following zero discount verstion of the optimal trajectory for undiscounted contributions:
$\log c_{t}^{*}[\infty]=\frac{\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}\right)$
The corresponding population value can hence be evaluated by the infinite sum of the undiscounted contributions, i.e.,
$\mathcal{V}[\infty]=\sum_{t=0}^{\infty}\log c_{t}^{*}[\infty]=$ $\frac{\infty\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)+\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}}{1-\theta}$
$=$ $\begin{cases}+\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta>1\\ {\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta=1\\ -\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}={A}\Theta<1\end{cases}$
The above result indicates that if ${A}\Theta>1$, then the population value is ever increasing and $n\to\infty$ must be the optimal solution; whereas, if ${A}\Theta<1$, then $n\to\infty$ must not be optimal and the population value must be maximized at a finite horizon $n^{*}\ll\infty$. The case where ${A}\Theta=1$ turns out to be the one that the planning horizon $n$ does not matter (beyond certain length) in maximizing the population value. Case V reflects ${A}\Theta>1$ where $n^{*}=\infty$, case VII reflects ${A}\Theta<1$ where $n^{*}\ll\infty$, and case VI reflects ${A}\Theta=1$ where $n^{*}$ is any number greater than $281$.888 The number is subject to decimals rounding. Figure 4 dipicts the population value function $\mathcal{V}[n]$, the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[n]$ and that of capital intensity $k_{t}^{*}[n]$, given the planning horizon $n$, for sample ZD settings V, VI, and VII.
The final case VIII implies ZD and AK setting ($\theta=\beta=1$), where $A\Theta=A=1+\rho>1$, under complete depreciation, by virtue of (13). Thus, if $\rho>0$ then $A>1$ in which case the population value would explode as described above, and we are left with $n^{*}\to\infty$. If in turn, $A=1+\rho=1$ as the case (VIII) goes, the infinite horizon population value is evaluated as follows:
$\mathcal{V}[\infty]=\lim_{\theta\to 1}{\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)=-\infty$
This indicates that there exists a finite optimum horizon $n^{*}\ll\infty$. Note that the population value function in this setting can be specified by zero discounting $(\beta\to 1)$ the AK model’s population value function (10), as follows:
$\mathcal{V}[n]$ $=\lim_{\beta\to 1}\left(\frac{\beta-((1-\beta){n}+1)\beta^{{n}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}{A}k_{0}\right)\right)$
$=\lim_{\beta\to 1}\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}k_{0}\right)=\lim_{\beta\to 1}(n+1)\beta^{n}\log\left(\frac{k_{0}}{(n+1)\beta^{n}}\right)$
$=(n+1)\log\left(\frac{k_{0}}{n+1}\right)$
The proof for the second identity, given $A=1$, is appended to Appendix 3. The third identity is subject to L’Hôpital’s rule. By the first order condition $\frac{\partial\mathcal{V}[n]}{\partial n}=0$, the optimal planning horizon is evaluated as $n^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
Figure 4: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right). Note that the population value function of the top-left panel is increasing indefinitely.
3.3 Intergenerational inequality
It seems obvious from Figure 3 (middle row) that consumption inequality among generations is worsened as we expand the planning horizon in the presence of future discounting (except for the knife-edge case III with $\log(A\beta)=0$), while it seems to improve in Figure 4 (middle row), for cases with zero discounting. To check the validity of the conjecture, we prepare Figure 5 to illustrate how Lorenz curve moves in regard to the planning horizon (i.e., $n=200,400,600$) for AK setting cases I, III, IV, and for ZD setting cases V, VI, and VII. The Lorenz curves (in Figure 5) are based on the stream of optimal consumption for a given planning horizon, namely, $c_{t}^{*}[n]$. The level of inequality increases in AK settings as the planning horizon is stretched longer. On the other hand, the level of inequality is relatively insensitive to the planning horizon for zero discounting cases, except for case VI with $\log(A\Theta)=0$ where inequality decreases along with the length of the planning horizon.
Figure 5: Lorenz curves for $n=200$ (solid line), $n=400$ (dotted line), and $n=600$ (dashed line). Top row panels (from left to right) correspond to AK settings I, III, and IV. Second row panels (from left to right) correspond to ZD settings V, VI, and VII.
For further analysis, let us consider Gini Index, a popular measure of inequality, defined on the optimal stream of consumptions $c_{t}^{*}[n]$, as follows:
$\mathcal{G}[n]=\frac{\sum_{t^{\prime}=0}^{n}\sum_{t=0}^{n}\left|c_{t}^{*}[n]-c_{t^{\prime}}^{*}[n]\right|}{2n\sum_{t=0}^{n}c_{t}^{*}[n]}$
Figure 6 illustrates Gini Index $\mathcal{G}$ as a function of planning horizon $n$ in various parameter settings. For both figures the the underlying parameters are fixed at $k_{0}=150$ and $A=1$. The left figure pertains to AK settings ($\theta=1$), while we raise the discount rate from $\beta=0.9$ (solid line) to $\beta=0.99$ (dashed line). As we can see from the (left) panel, AK settings have the tendency to unequalize consumptions between generations as the planning horizon expands, while inequality is mitigated by less discountig the future. In contrast, the right figure indicates that ZD settings have the tendency to equalize consumptions between generations as the planning horizon expands, while inequality is enhanced by the higher output elasticity of capital.
Figure 6: The panels are planning horizon vs Gini Index i.e., $(n,\mathcal{G}[n])$ for AK model cases with different discount rates (left), and zero discounting cases with different output elasticities of capital (right). The parameters setting of the left panel is $\theta=1$, $k_{0}=150$, $A=1$, $\beta=0.9$ (solid line), and $\beta=0.99$ (dashed line). The parameters setting of the right panel is $\beta=1$, $k_{0}=150$, $A=1$, $\theta=0.9$ (solid line), and $\theta=0.95$ (dashed line).
3.4 The initial action
Let us now focus on the initial action of the optimal consumption schedule which we specify, by virtue of (7) or (21), as follows:
$c_{0}^{*}[n]=\frac{{A}(k_{0})^{\theta}}{S_{n}}=\frac{1-\beta\theta}{1-(\beta\theta)^{n}}{A}(k_{0})^{\theta}$ (18)
We take the derivative of (18) and obtain the following result:
$\frac{\mathrm{d}c_{0}^{*}[n]}{\mathrm{d}n}=\frac{A(k_{0})^{\theta}\left(1-\beta\theta\right)(\beta\theta)^{n}\log(\beta\theta)}{\left(1-(\beta\theta)^{n}\right)^{2}}<0$
since, $\beta\theta<1$. In other words, the further we take future generations into account, the more current generation must sacrifice. In any case, the initial action is monotonously decreasing in the planning horizon, and ultimately converge to the following:
$c_{0}^{*}[\infty]=(1-\beta\theta){A}(k_{0})^{\theta}$ (19)
By taking it’s derivative with respect to $\beta$ we obtain:
$\frac{\partial c_{0}^{*}[\infty]}{\partial\beta}=-\theta A(k_{0})^{\theta}<0$
That is to say, the more we respect future generations (by raising $\beta$), the more current generation must sacrefice. From another perspective, we can solve the equation $c_{0}^{*}[n]=s$ for $n$ in regard to (18), where $s$ is the consumption level of subsistence, as follows:
$n=\frac{\log\left((1-\beta\theta)A(k_{0})^{\theta}\right)-\log s}{\log(\beta\theta)}$
This $n$ is the subsistence-proof number of potential generations.
Finally, for AK models, $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$ by (19), but this value will approach zero if future generations are respected as equally as possible to the current one ($\beta\to 1$). That is to say, the ultimate respect for the future generations in AK models with an infinite planning horizon induces the current generation to sacrefice ultimately, i.e., $c_{0}^{*}[\infty]\to{s}$. Alternatively, if we relax the infinite horizon assumption, initial action for AK models with ultimate respect for the future generations be:
$c_{0}^{*}[n]=\lim_{\beta\to 1}\frac{1-\beta}{1-\beta^{n}}{A}k_{0}=\frac{Ak_{0}}{n}$
and we are left with the AK and ZD version of subsistence-proof number of potential generations $n=Ak_{0}/s$.
4 Concluding Remarks
The extent that we take into account of the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor. Notice that the potential number of intrinsic years that are accounted by compound discounting at a rate of $r$ is as much as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$. As regards to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting at a planning horizon $n=1/r$. That is, for example, a generational discount rate of $1$% amounts to 100 generations. Binary discounting is flexible in a sense that the planning horizon is variable, so that infinity is not precluded from an option. Moreover, it is nondiscriminatory, ensuring that living generations are treated equally. With binary discounting, however, an economic growth model requires a finite horizon dynamic optimization approach.
One important feature of binary discounting is that it hinders future generations that earn below-critical-level utilities. If the phase boundary $\log(A\beta)$ for AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level), is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of subsistence. Binary discounting avoids such repugnancy by limiting the longevity of the dynasty. On the other hand, if the phase boundary $\log({A}\Theta)$ for zero (compound) discounting is negative, a finite horizon must be optimal since an infinite horizon objective function approaches a negative infinity. Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4), yield increasingly negative contributions to the population value.
As regards the distribution of consumption among generations, the rate of compound discounting plays the key role. That is, the less the rate of discounting is, the more equal the consumptions become. Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in the AK setting cases, whereas vice-versa in a zero discounting setting. Our another concern was the optimal initial consumption in light of the well-being of future generations. In any case the initial consumption is decreasing in the planning horizon, and ultimately, it will depend on the value of $\beta\theta$. The optimal initial consumption for infinite planning horizon will decrease as discount rate approaches zero, and under the AK setting, initial consumption itself approaches zero until it reaches the level of subsistence.
Appendix 1
Below we put down a production function and its dual unit cost function:
${Y}=A{K}^{\theta}{L}^{1-\theta},$ $p={B}^{-1}{r}^{\theta}w^{1-\theta}$
where $p$, $r$, and $w$ denote prices for $Y$, $K$, and $L$, respectively. The remaining parameter $A$ is called the productivity and $B$ is called the dual cost function-based productivity. Applying Shephard’s lemma to the dual function leads to the followings.
$\frac{\partial p}{\partial r}=\frac{\theta}{B}\left(\frac{r}{w}\right)^{\theta-1}=\frac{K}{Y},$ $\frac{\partial p}{\partial w}=\frac{1-\theta}{B}\left(\frac{r}{w}\right)^{\theta}=\frac{L}{Y}$
By virtue of these euqations, the marginal product of capital (MPK) can be readily evaluated as follows:
$\text{MPK}=\frac{\partial Y}{\partial K}=A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=A\theta^{\theta}(1-\theta)^{1-\theta}\left(\frac{r}{w}\right)^{\theta-1}$
$A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=\frac{Y}{K}\theta=B\left(\frac{r}{w}\right)^{\theta-1}$
where, $r/w\in(0,\infty)$ equals the marginal rate of substitution between capital and labor (MRS). By comparison, we are left with $B=A\theta^{\theta}(1-\theta)^{1-\theta}\equiv A\Theta$. Moreover,
$A=\left(\frac{\text{MPK}}{\Theta}\right)\left(\frac{r}{w}\right)^{\theta-1}$ $B=\text{MPK}\left(\frac{r}{w}\right)^{\theta-1}$
On the other hand, recall the breakdown equation of total output (2) and take the partial derivative as follows:
$\frac{\partial Y_{t}}{\partial K_{t}}=\left(\frac{\partial K_{t+1}}{\partial K_{t}}-1\right)+\delta=\rho+\delta=\text{MPK}$ (20)
where $\rho>0$ denotes the rate of interest. Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$.
According to the study in section 3.1, we note that condition $A\beta<1$ implies the non-increasing of undiscounted contributions (or utilities less the critical level) with respect to generations. The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations. Figure 7 (left) dipicts the set of possible values for $(\theta,A)$ that satisfies $A\beta\leq 1$. Also, according to the study in section 3.2, we note that condition $A\Theta<1$ implies the existence of a finite optimal planning horizon. Otherwise, if $A\Theta>1$, the optimal planning horizon will be inifinite. The knife-edge case where $A\Theta=1$ implies that the optimal planning horizon is indeterminate. Figure 7 (right) dipicts the set of possible values for $(\theta,A)$ that satisfies $A\Theta\leq 1$.
[…]
Figure 7: The shaded area (mariginal lines not included) corresponds to the set of parameters $(\theta,A)$ where $B=A\theta^{\theta}(1-\theta)^{1-\theta}<1$
Appendix 2
Lemma 1.
The value function of Bellman equation (4) is of the following:
$\mathcal{V}_{t}\left[k_{t};n\right]$ $=S_{{n}-t}\theta\log(k_{t})+{R}_{t}$
where $S_{n-t}$ is specified by (6) and ${R}_{t}$ is a term that does not depend on $k_{t}$.
Proof.
We show this by induction. First, evaluate the value function at $t={n}$ by (4):
$\mathcal{V}_{n}\left[k_{n};{n}\right]=\max_{c_{n}}\left(\log c_{n}+\beta\mathcal{V}_{n+1}[k_{{n}+1}={A}(k_{n})^{\theta}-c_{n};{n}]\right)$
Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Also, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, we know that the final maximization is bounded, i.e.,
$\mathcal{V}_{n}\left[k_{n};{n}\right]=\log c_{n}=\log{A}+\theta\log k_{n}$
Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the proposition holds true for $t={n}$.
Suppose the proposition holds true for $t+1$. Then,
$\mathcal{V}_{t}\left[k_{t};n\right]=$ $\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{V}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};n\right]\right)$
$=$ $\max_{c_{t}}\left(\log c_{t}+\beta\left(S_{{n}-t-1}\theta\log\left({A}(k_{t})^{\theta}-c_{t}\right)+{R}_{t+1}\right)\right)$
Below are the corresponding first order condition and its solution:
$\frac{1}{c_{t}}-\frac{(\beta\theta)S_{{n}-t-1}}{{A}(k_{t})^{\theta}-c_{t}}=0,$ or, $c_{t}=\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}$ (21)
Here, we used $(\beta\theta)S_{{n}-t-1}=\beta\theta+(\beta\theta)^{2}+\cdots+(\beta\theta)^{n-t}=S_{n-t}-1$. By plugging the above solution back into the maximand we arrive at the following result:
$\mathcal{V}_{t}[k_{t};n]$ $=\log\left(\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}\right)+\beta\left(R_{t+1}+S_{n-t-1}\theta\log\left({A}(k_{t})^{\theta}-\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}\right)\right)$
$=S_{{n}-t}\theta\log(k_{t})+\left(\beta{R}_{t+1}+S_{{n}-1}\log\left(\frac{{A}}{S_{{n}-t}}\right)+\log(S_{n-1}-1)^{S_{n-1}-1}\right)$
$=S_{{n}-t}\theta\log(k_{t})+R_{t}$
Hence, the lemma. ∎
Lemma 2.
The optimal trajectory of state $k_{t}^{*}[n]$ for Bellman equation (4) is:
$\ln k_{t}^{*}[n]=\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}$
Proof.
We show this by induction. By plugging (21) into (2.1), we obtain:
$k_{t+1}={A}(k_{t})^{\theta}-c_{t}=\frac{S_{{n}-t}-1}{S_{{n}-t}}{A}(k_{t})^{\theta}=\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta(k_{t})^{\theta}$ (22)
As we apply $t=0$ on the above (22) and by taking the logarithm,
$\log k_{1}=\ln\left(\frac{S_{{n}-1}}{S_{n}}{A}\beta\theta\right)+\theta\log k_{0}$
we know that the proposition is true for $t=1$.
Suppose that the proposition is true for $t$. Then, we know by (22) that:
$\log k_{t+1}^{*}[n]$ $=\log\left(\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta\right)+\theta\log k_{t}^{*}[n]$
$=\log\left(\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
$=\sum_{i=1}^{t+1}\theta^{t+1-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t+1}\log k_{0}$
That is, the proposition is true for $t+1$. ∎
Proposition 1.
The optimal consumption trajectory $c_{t}^{*}[n]$ for Bellman equation (4) is:
$\log c_{t}^{*}[n]=\log\left(\frac{{A}}{S_{{n}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
Proof.
Obvious from Proposition 2 and (21). ∎
Appendix 3
Let us evaluate the following:
$\lim_{\beta\to 1}\frac{\beta-((1-\beta){n}+1)\beta^{n+1}}{(1-\beta)^{2}}$ $=\lim_{\beta\to 1}\frac{1-(n+1)((1-\beta)n+1)\beta^{n}+n\beta^{n+1}}{-2(1-\beta)}$
$=\lim_{\beta\to 1}\frac{n(n+1)\left(2\beta^{n}-((1-\beta)n+1)\beta^{n-1}\right)}{2}$
$={n(n+1)}/{2}$
where we used the L’Hôpital’s rule twice. Then, we know that
$\lim_{\beta\to 1}\frac{\beta-((1-\beta){n}+1)\beta^{n+1}}{(1-\beta)^{2}}\log(A\beta)=\frac{n(n+1)}{2}\log(A)=0$