### v3

Abstract
This paper develops a dynamic optimization framework for evaluating the optimal longevity of a dynasty under a critical-level utilitarian social welfare criterion. Unlike standard models that assume an infinite planning horizon and apply discounting to future generations, we endogenize the planning horizon itself and examine the welfare-maximizing number of generations in a roundabout production economy with capital accumulation. The model incorporates a well-being threshold that must be met for generational existence to be considered ethically justified, providing an alternative to traditional utilitarian aggregation. Using finite-horizon dynamic programming, we derive closed-form consumption trajectories and analyze how technological parameters and intertemporal preferences affect the optimal generational span. Our results show that the optimal horizon need not be infinite—even under zero discounting—and that intergenerational equity may require limiting dynasty length.
[…]
O21 I30 C61
Keywords:
Critical-level Utilitarianism Dynamic Programming Optimal Population Intergenerational Equity
journal:
1 Introduction
Economic models of intertemporal planning often assume an infinite horizon, implicitly treating dynasties or economies as perpetually forward-looking entities. However, such assumptions can obscure important trade-offs between current and future welfare, particularly when the viability of future generations depends on limited resources and capital dynamics. In this paper, we examine a social planner’s problem where the planning horizon—the number of generations considered—is endogenously chosen. By adopting a critical-level utilitarian welfare function, we evaluate the contribution of each generation relative to a minimum well-being threshold. This framework allows us to study how productivity, capital accumulation, and time preference interact to determine the ethically and economically optimal duration of a dynasty.
We begin by positing a (classical) unitilitarian social planner whose objective is to indentify a state $x$ that maximizes the sum of the utilities $\mathcal{V}=\sum_{i\in n(x)}\Upsilon_{i}(x)$, which we refer to as the population value, where $n(x)$ denotes the set of individuals and $\Upsilon_{i}(x)$ denotes the lifetime utility (i.e., the level of well-being) of an individual $i$. For simplicity, let us assume that all individuals are identical and consume an equal share, denoted by $c$, from a fixed amount of resource $R>0$. A utilitarian axiology examines the problem of finding the population size $n=R/c$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\Upsilon(c_{i})=n\Upsilon(c=R/n)$. The utility function $\Upsilon(c)$ is assumed to be monotonically increasing and concave. In addition, with respect to the population value $\mathcal{V}=n\Upsilon(c)$, zero utility ($\Upsilon=0$) is equivalent to the absence of life ($n=0$).
The consumption level $c=\nu$, where $\Upsilon(\nu)=0$, be referred to as the well-being subsistence level, following Dasgupta (2019). Owing to the monotonicity of the utility function, $\Upsilon(c)>0$ must hold as long as $c>\nu$ (see Figure 1 (left)). It then follows that the domain allowed for $c$ is $(\nu,\infty)$ since $\Upsilon(c)$ represents the level of attainable lifetime utility. Figure 1 (left) depicts the optimal solution $c=\omega$, which satisfies the first-order condition for maximizing the population value, i.e., $\Upsilon^{\prime}(\omega)=\Upsilon(\omega)/\omega$. Owing to the concavity of the utility function, if the well-being subsistence approaches zero ($\nu\to 0$), as Córdoba (2023) refers to the homothetic case, the optimal consumption also approaches zero ($\omega\to\nu$), while the optimal size of the population approaches infinity ($n^{*}=R/\omega\to\infty$), leading to a conclusion that is considered repugnant (Parfit, 1986).
Meanwhile, Blackorby et al (2005) introduced a population axiology known as critical-level utilitarianism. Within this framework, a social planner aims to determine a set of consumption schedules $c_{i}$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)$, where an individual’s contribution to the population value is evaluated relative to the critical utility level $\alpha$ (see Figure 1 (right)). That is, an individual $i$’s life contributes positively to the population value if $c_{i}>\kappa=\Upsilon^{-1}(\alpha)$ but negatively otherwise. Note that setting $\alpha>0$ prevents the repugnant conclusion even when the underlying utility function is homothetic ($\nu=0$). For convenience, we employ a utility function of the form $\Upsilon(c_{i})=\log c_{i}+\alpha$ for analyses so that the value function for the critical-level utilitarian population value in our maximization study becomes $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)=\sum_{i=1}^{n}\log c_{i}$.
As we recall our primary idea of incorporating roundabout production into the optimal population framework, let us introduce (discrete) time and assume that the number of individuals per period is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ during which they live. [^1] We may therefore refer to $t$ as a generation. The population value function then becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$. Note that this type of value function was widely employed in the previous Ramsey–Cass–Koopmans type of dynastic optimization model (including Nordhaus, 1992; Stern, 2007), except that the future generations were discounted according to the temporal deviation from the present, and typically, the number of potential generations $n$ remains unquestioned (and set to infinity). In contrast, in this study, potential generations are accounted for equally, and the number of generations is optimized.
The remainder of the paper proceeds as follows. In the following section, we first describe the model of roundabout production on the basis of Cobb–Douglas technology. We then define our main problem in the form of finite horizon dynamic programming, where the objective function comprises a series of Bernoullian contributions, and analytically derive a general solution. In Section 3, we introduce two broad settings of parameters to simplify the general solution and allow us to study the optimal trajectory of the variables in a numerical fashion. The first setting is the AK setting, where the output elasticity of capital is set to unity. The second is the zero discounting (ZD) setting, where the discount factor is set to unity. We also study the peculiarity of the two models with respect to intergenerational equity and the optimal initial action of consumption. Section 4 concludes the paper.
Figure 1: Left: Utility function of an individual. The consumption level for neutrality (or well-being subsistence) is denoted by $\nu$. Marginal utility and average utility coincide at $c=\omega$. Right: Contribution to the critical-level utilitarian population value, where $\alpha$ denotes the critical level. At $c=\kappa$, the utility reaches the critical level. In all cases, the domain of permissible consumption levels $c$ for existence is $c>\nu$.
2 Model
2.1 Roundabout production
Let us begin by postulating a two-factor Cobb–Douglas production function characterized by constant returns to scale, as follows:
$Y_{t}={A}(K_{t})^{\theta}(L_{t})^{1-\theta}$ (1)
where $Y_{t}$ denotes the economy’s output, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all of which are effective during period $t$. For the relevant parameters, $10<\theta<$ 1 denotes the output elasticity of capital. The level of technology is indicated by productivity, denoted by $A$, which is assumed to be fixed throughout the considered time span. Note that ${B}\equiv{A}\theta^{\theta}(1-\theta)^{1-\theta}$ is the cost function-based productivity. [^2] Otherwise, $B$ is referred to as the dual productivity. See Appendix 1 for more details.
The breakdown of output into investment, capital depreciation, and final consumption is described below:
$Y_{t}=K_{t+1}-K_{t}+\delta K_{t}+C_{t}$ (2)
Here, $\delta\leq 1$ denotes the capital depreciation factor, and $C_{t}$ denotes aggregate consumption in period $t$. Combining equations (1) and (2) and dividing both sides by $L_{t}=L$, which we assume to be constant over time, leads to the following intertemporal dynamics (or roundaboutness) of capital intensity:
$y_{t}={A}(k_{t})^{\theta}=k_{t+1}-(1-\delta)k_{t}+c_{t}$ (3)
where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in periods $t$ and $t+1$, respectively. Additionally, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per capita output and consumption, respectively, in period $t$.
[…]
2.2 Finite horizon dynamic programming
The primary problem here is to solve for an optimal consumption path given a planning horizon. The Bellman equation of the problem is as follows:444 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
$\mathcal{\mathcal{V}}_{t}\left[k_{t};\ n\right]=\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{\mathcal{V}}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};\ n\right]\right)$ (4)
The optimal trajectory of the state variable $k_{t}$ is given by Lemma 2, which we append to Appendix 2 with proofs, as follows:
$k_{t}^{*}[n]=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)^{\theta^{t-i}}=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i-1}}$ (5)
where $S_{\ell}$ is defined as follows:
$S_{\ell}\equiv\sum_{i=0}^{\ell}(\beta\theta)^{i}=1+(\beta\theta)+(\beta\theta)^{2}+\cdots+(\beta\theta)^{\ell}=\frac{1-(\beta\theta)^{\ell+1}}{1-\beta\theta}$ (6)
The following optimal trajectory of consumption is obtained by (5) and (21) Appendix 2, which must be true according to the proof of Lemma 1.
$c_{t}^{*}[n]=\frac{{A}\left(k_{t}^{*}[n]\right)^{\theta}}{S_{n-1}}=\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{n}-t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i}}$ (7)
With formula (7), the population value function is given as follows:
$\mathcal{V}[n]$ $=\sum_{t=0}^{n}\beta^{t}\log c_{t}^{*}[n]$
$=\log\left(\frac{{A}(k_{0})^{\theta}}{S_{n}}\right)+\sum_{t=1}^{n}\beta^{t}\log\left(\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{n}-t}}\prod_{i=1}^{t}\left(\frac{S_{{n}-t+i-1}}{S_{{n}-t+i}}{A}\beta\theta\right)^{\theta^{i}}\right)$ (8)
We hereafter aim to maximize this function with respect to the planning horizon $n$.
Our approach to analyzing the population value (8) adopts a numerical rather than an analytical framework because the derivative of $\mathcal{V}[n]$ with respect to $n$ does not seem to provide meaningful insights. In the following section, the optimal trajectory of contributions $\log c_{t}^{*}[n]$ and the population value $\mathcal{V}[n]$, for any given planning horizon $n$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$. We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon is evaluable under $\beta\theta<1$. We therefore base our study of ZD ($\beta=1$) on this parameter setting (i.e., $\beta\theta<1$, which indicates that $\theta<1$, consistent with a Cobb–Douglas model). In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$), which we term AK settings, and Cobb–Douglas production without future discounting ($\theta<1$ and $\beta=1$), which we term ZD settings.
Table 1 summarizes the parameter settings chosen for the numerical examinations. Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log(A\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log B$. Cases VIII and IX correspond to the parameter settings where $\log({A}\beta)=\log B=0$.
Table 1: Parameters applied in various cases.
case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log B$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.7$
II $1.01$ $0.992$ $1$ $+$ $95$ $60.0$
III $1/\beta$ $0.992$ $1$ 00 $73$ $55.6$
IV $1.005$ $0.992$ $1$- $58$ $51.0$
V $1.05$ $1$ $0.992$ $+$ $\infty$ +$\infty$+ ∞
VI $1.2$ $1$1 00 $(281)$ $41.7$
VII $1.05$ $1$ $0.991$- $117$$\infty$- ∞
VIII $1$ $1$ $1$ 00 00 $54$ $55.2$$\infty$- ∞
IX $1/\beta$ $0.992$2 00 00 $53$ $49.1$ $-1843.2$
  • Note:
    For all cases I–IX, $k_{0}=150$. $B=A(1-\theta)^{1-\theta}\theta^{\theta}$ is the cost-function based productivity.
  • *1
    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=B=$ 1.
  • *2
    $\theta\approx 0.998982$ where $(1/0.992)(1-\theta)^{1-\theta}\theta^{\theta}=B=1$.
3 Analysis
3.1 AK setting
The AK model, which was formally developed by Frankel (1962), is one of the simplest fundamental models of endogenous growth. Here, we employ this production model to study the population value function that discounts future generations to determine whether $n\to\infty$ is an optimal policy. The optimal consumption path (7) for the AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
$c_{t}^{*}[n]=\left(\frac{{A}k_{0}}{S_{{n}-t}}\right)\left(\frac{S_{{n}-1}}{S_{n}}{A}\beta\right)\left(\frac{S_{{n}-2}}{S_{{n}-1}}{A}\beta\right)\cdots\left(\frac{S_{{n}-t}}{S_{{n}-t+1}}{A}\beta\right)=\frac{({A}\beta)^{t}{A}k_{0}}{S_{n}}$ (9)
The population value function (8), therefore, becomes:
$\mathcal{V}[n]$ $=\sum_{t=0}^{n}\beta^{t}\log\left(\frac{({A}\beta)^{t}{A}k_{0}}{S_{n}=1+\beta+\cdots+\beta^{n}}\right)$
$=\frac{\beta-((1-\beta){n}+1)\beta^{{n}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}{A}k_{0}\right)$ (10)
For the sake of the analysis, let us take the derivative with respect to ${n}$.
$\frac{\mathrm{d}\mathcal{V}[n]}{\mathrm{d}n}=\lambda\left(\gamma-{n}\log({A}\beta)+\log\left(1-\beta^{n+1}\right)\right)$ (11)
where $\lambda$ and $\gamma$ are given as follows:
$\lambda=\frac{\beta^{n}\log\beta^{\beta}}{1-\beta}>0,$ $\gamma=1-\frac{\log({A}\beta)^{1-\beta+\log\beta}}{\log\beta^{1-\beta}}-\log({A}(1-\beta)k_{0})$
To study the derivative sign of (11), we consider the following two functions, whose equal values, i.e., $f[n]=g[n]$, convey the first-order condition of optimality.
$f[n]=-\gamma+n\log({A}\beta),$ $g[n]=\log\left(1-\beta^{n+1}\right)$ (12)
Before we proceed, let us examine the possible range of the parameter $A$. The evaluation of the marginal product of capital (MPK) of our production function (1) net of depreciation, which should coincide with the real interest rate $\rho>0$, viz.,
$\frac{\partial Y_{t}}{\partial K_{t}}-\delta={A}\theta\left(\frac{K_{t}}{L_{t}}\right)^{\theta-1}-\delta=\rho>0$ (13)
In AK models with complete capital depreciation, $\theta=\delta=1$ leads to ${A}=1+\rho$. Therefore, assuming that ${A}>1$ is relevant in this setting.555 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross interest rate and generational preference rate of the population.
Figure 2 (left) depicts the functions $f[n]$ and $g[n]$ under different parameters for cases I–IV. Clearly, $f[n]$ is a linear function whose slope is $\log({A}\beta)$, and $g[n]$ monotonously increases and approaches zero, i.e., $g[\infty]=0$. Here, we fix the discount factor $\beta$ and differentiate the productivity ${A}$ at four different levels. As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[n]$ will intersect with $g[n]$ at a single point, and the population value function $\mathcal{V}[n]$ has a single peak. If ${A}>1/\beta$ so that $f[n]$ has a positive slope, then $f[n]$ and $g[n]$ could either intersect with two points where the population value $\mathcal{V}[n]$ may rise, fall and then rise again, or never intersect so that the population value may rise indefinitely with respect to the planning horizon $n$. Figure 2 (right) depicts the population value functions for cases I–IV.
Figure 2: Left: The solid lines represent $g[n]$, as defined in (12), for cases I–IV with clear correspondences, i.e., the steepest slope corresponds to case I, while case IV corresponds to the scenario with a negative slope. The dashed line represents $f[n]$, which is also defined in (12). Right: The plots show population value functions $\mathcal{V}[n]$ for cases I–IV with obvious correspondences.
To visualize the optimal trajectory of the variables in various situations, we specify them here in the form of functionals. By referencing (9), the optimal trajectory of the undiscounted contribution to the population value becomes linear with respect to $t$, as follows:
$\log c_{t}^{*}[n]=\log\left(\frac{1-\beta}{1-\beta^{n+1}}{A}k_{0}\right)+t\log({A}\beta)$ (14)
In reference to (21) Appendix 2, the optimal trajectory of capital intensity becomes as follows:
$k_{t}^{*}[n]=\frac{S_{n-t}}{{A}}c_{t}^{*}[n]=\frac{1-\beta^{n+1-t}}{1-\beta^{n+1}}({A}\beta)^{t}k_{0}$ (15)
Figure 3 displays the population value function $\mathcal{V}[n]$, specified as (10), for cases I, III, and IV on the top. [^6] These figures correspond to those depicted in Figure 2 (right). Note that, as long as $\beta<1$, the population value function will always converge to a finite value, regardless of the parameter settings in addition to $\beta$. As we let $n\to\infty$ in (10), we have the following:
$\mathcal{V}[\infty]=\frac{\log\left({A}(1-\beta)^{1-\beta}\beta^{\beta}\right)}{(1-\beta)^{2}}+\frac{\log(k_{0})}{1-\beta}$ (16)
By comparing the population values obtained at the numerical solution of the first-order condition $f[n]=g[n]$, as referenced in (12), with $\mathcal{V}[\infty]$ from (16), we know the optimal planning horizon $n^{*}$, which we display in Table 1 for cases I–IV. The middle row of Figure 3 displays the optimal trajectories of the undiscounted contributions to the population value, as described in (14), for planning horizons $n=200,400,600$ for cases I, III, and IV (from left to right). Similarly, the bottom row of Figure 3 displays the optimal trajectories of capital intensity, based on (15), for planning horizons $n=200,400,600$ for cases I, III, and IV (from left to right).
Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK setting cases I (left), III (center), and IV (right). The parameters corresponding to each of the cases are given in Table 1.
3.2 ZD setting
Here, we study ZD, i.e., $\beta=1$, under Cobb–Douglas production with an ordinary output elasticity of capital $\theta<1$. The condition that $\beta\theta<1$ nonetheless allows us to evaluate the key summation as follows:
$S_{\infty-\tau}=\lim_{n\to\infty}\frac{1-(\beta\theta)^{n-\tau+1}}{1-\beta\theta}=\frac{1}{1-\beta\theta}$
We apply the above and $n\to\infty$ to (7) and obtain the following:
$\log c_{t}^{*}[\infty]$ $=\log\left(\frac{{A}}{S_{{\infty}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{\infty}-i}}{S_{{\infty}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
$=\log\left({{A}}{(1-\beta\theta)}\right)+\left(\theta^{t-1}+\cdots+\theta+1\right)\log({A}\beta\theta)^{\theta}+\theta^{t+1}\log k_{0}$
$=\frac{\log\left({A}(1-\beta\theta)^{1-\theta}(\beta\theta)^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\beta\theta)^{\frac{\theta}{1-\theta}}\right)$ (17)
where the optimal trajectory for undiscounted contributions is geometrically convergent. We can then apply $\beta=1$ to arrive at the following zero discount verification of the optimal trajectory for undiscounted contributions:
$\log c_{t}^{*}[\infty]=\frac{\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}\right)$
The corresponding population value can hence be evaluated by the infinite sum of the undiscounted contributions, i.e.,
$\mathcal{V}[\infty]=\sum_{t=0}^{\infty}\log c_{t}^{*}[\infty]=$ $\frac{\infty\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)+\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}}{1-\theta}$
$=$ $\begin{cases}+\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B>1\\ {\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B=1\\ -\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B<1\end{cases}$
The above result indicates that if $B>1$, then the population value is ever increasing, and $n\to\infty$ must be the optimal solution. However, if $B<1$, then $n\to\infty$ must not be optimal, and the population value must be maximized at a finite horizon $n^{*}\ll\infty$. The case where $B=1$ is the knife-edge case in which the planning horizon $n$ does not matter (beyond a certain length) in maximizing the population value. Case V corresponds to $B>1$, and $n^{*}=\infty$; case VII corresponds to $B<1$, and $n^{*}\ll\infty$; and case VI corresponds to $B=1$, and $n^{*}$ is any number greater than $281$.777 The number is subject to decimal rounding. Figure 4 depicts the population value function $\mathcal{V}[n]$, along with the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[n]$ and the capital intensity $k_{t}^{*}[n]$, given a planning horizon $n$, for sample ZD settings V, VI, and VII.
The final case VIII relates to the ZD/AK setting ($\theta=\beta=1$), where $B=A=1+\rho>1$, under complete depreciation, as described in (13). Thus, if $\rho>0$, then $A>1$, in which case the population value would increase without bound, as described above, leading to $n^{*}\to\infty$. If, in turn, $A=1+\rho=1$, as in case VIII, the infinite horizon population value is evaluated as follows:
$\mathcal{V}[\infty]=\lim_{\theta\to 1}{\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)=-\infty$
This indicates that there exists a finite optimum horizon $n^{*}\ll\infty$. Note that the population value function in this setting can be specified by applying a unitary discounting factor $(\beta\to 1)$ to the population value function (10) of AK models as follows:
$\mathcal{V}[n]$ $=\lim_{\beta\to 1}\left(\frac{\beta-((1-\beta){n}+1)\beta^{{n}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}{A}k_{0}\right)\right)$
$=\lim_{\beta\to 1}\frac{1-\beta^{{n}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{n}+1}}k_{0}\right)=\lim_{\beta\to 1}(n+1)\beta^{n}\log\left(\frac{k_{0}}{(n+1)\beta^{n}}\right)$
$=(n+1)\log\left(\frac{k_{0}}{n+1}\right)$
The proof for the second identity, given $A=1$, is appended to Appendix 3. The third identity is subject to L’Hôpital’s rule. By the first-order condition $\frac{\partial\mathcal{V}[n]}{\partial n}=0$, the optimal planning horizon is evaluated as $n^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
Figure 4: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right). Note that the population value function of the top-left panel is increasing indefinitely. The parameters corresponding to each of the cases are given in Table 1.
3.3 Intergenerational inequality
Figure 3 (middle row) shows that consumption inequality across generations increases as the planning horizon extends under future discounting (except in the knife-edge case III when $\log(A\beta)=0$), whereas it seems to decrease in Figure 4 (middle row) for cases with ZD. To validate this conjecture, Figure 5 shows how the Lorenz curve shifts with respect to the planning horizon (namely, $n=200,400,600$) for AK setting cases I, III, and IV and ZD setting cases V, VI, and VII. These Lorenz curves (of Figure 5) are based on the sequence of optimal consumption levels for a given planning horizon, i.e., $c_{t}^{*}[n]$. The inequality level increases in AK settings as the planning horizon expands. On the other hand, inequality level is relatively insensitive to the planning horizon for ZD setting cases, except for case VI, when $\log B=0$, where inequality decreases as the planning horizon increases.
Figure 5: Lorenz curves for $n=200$ (solid line), $n=400$ (dotted line), and $n=600$ (dashed line). The top row panels (from left to right) correspond to AK settings I, III, and IV. The second row panels (from left to right) correspond to ZD settings V, VI, and VII.
For further analysis, let us consider the Gini index, a popular measure of inequality, defined on the basis of the optimal stream of consumption $c_{t}^{*}[n]$, as follows:
$\mathcal{G}[n]=\frac{\sum_{t^{\prime}=0}^{n}\sum_{t=0}^{n}\left|c_{t}^{*}[n]-c_{t^{\prime}}^{*}[n]\right|}{2n\sum_{t=0}^{n}c_{t}^{*}[n]}$
Figure 6 shows the Gini index $\mathcal{G}$ as a function of the planning horizon $n$ under various parameter settings. For both panels, the underlying parameters are fixed at $k_{0}=150$ and $A=1$. The left panel corresponds to the AK settings ($\theta=1$) as the discount rate is increased from $\beta=0.9$ (solid line) to $\beta=0.99$ (dashed line). In the AK settings, consumption tends to become more unequal across generations as the planning horizon expands. However, this effect is mitigated when future discounting decreases. In contrast, the right panel indicates that ZD settings tend to equalize consumption across generations as the planning horizon expands, whereas inequality is enhanced by the higher output elasticity of capital.
Figure 6: The panels depict the planning horizon vs Gini Index, i.e., $(n,\mathcal{G}[n])$ for AK setting cases with different discount rates (left), and ZD setting cases with different output elasticities of capital (right). The parameter settings for the left panel are $\theta=1$, $k_{0}=150$, $A=1$, $\beta=0.9$ (solid line), and $\beta=0.99$ (dashed line). The parameter settings for the right panel are $\beta=1$, $k_{0}=150$, $A=1$, $\theta=0.9$ (solid line), and $\theta=0.95$ (dashed line).
3.4 Initial action
Let us focus on the initial action of the optimal consumption schedule, which we specify as follows based on (7) or (21):
$c_{0}^{*}[n]=\frac{{A}(k_{0})^{\theta}}{S_{n}}=\frac{1-\beta\theta}{1-(\beta\theta)^{n}}{A}(k_{0})^{\theta}$ (18)
We take the derivative of (18) and obtain the following result:
$\frac{\mathrm{d}c_{0}^{*}[n]}{\mathrm{d}n}=\frac{A(k_{0})^{\theta}\left(1-\beta\theta\right)(\beta\theta)^{n}\log(\beta\theta)}{\left(1-(\beta\theta)^{n}\right)^{2}}<0$
since $\beta\theta<1$. In other words, the more future generations are treated equally, the more current generation must reduce their consumption. In any case, the initial action monotonically decreases in the planning horizon and ultimately converges to the following:
$c_{0}^{*}[\infty]=(1-\beta\theta){A}(k_{0})^{\theta}$ (19)
By taking its derivative with respect to $\beta$, we have:
$\frac{\partial c_{0}^{*}[\infty]}{\partial\beta}=-\theta A(k_{0})^{\theta}<0$
In other words, the more we consider future generations (by raising $\beta$), the more the current generation must reduce their consumption. From another perspective, we can solve equation $c_{0}^{*}[n]=\nu$ for $n$ with respect to (18), where $\nu$ is the well-being subsistence, as follows:
$n=\frac{\log\left((1-\beta\theta)A(k_{0})^{\theta}\right)-\log\nu}{\log(\beta\theta)}$
This $n$ is the subsistence-proof size of potential generations.
Finally, for AK models, the optimal initial action by (19) is $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$, but this value approaches zero if future generations are given equal consideration as the current generation ($\beta\to 1$). That is, in AK models with an infinite planning horizon, giving ultimate consideration to future generations requires the current generation to ultimately reduce their consumption, i.e., $c_{0}^{*}[\infty]\to{\nu}$. Alternatively, if we relax the infinite horizon assumption, the initial action for AK models with ultimate consideration for the future generations becomes:
$c_{0}^{*}[n]=\lim_{\beta\to 1}\frac{1-\beta}{1-\beta^{n}}{A}k_{0}=\frac{Ak_{0}}{n}$
and we are left with the AK/ZD version of subsistence-proof size of potential generations, that is, $n=Ak_{0}/\nu$.
4 Concluding Remarks
The extent to which we account for the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor. The potential number of intrinsic years that are accounted for by compound discounting at a rate of $r$ is as high as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$. With respect to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting (which eliminates consideration for utilities before the planning horizon while fully discounting those that lie beyond it) at a planning horizon $n=1/r$. Binary discounting is flexible in the sense that the planning horizon is variable so that infinity is not precluded as an option. Moreover, it is nondiscriminatory, ensuring that existing generations are treated in ultimate equity. With binary discounting, however, the roundaboutness of production requires a finite horizon dynamic programming approach.
One important feature of our welfare evaluation on the basis of critical-level utilitarianism is that it hinders future generations that earn below-critical-level utility. If the phase boundary $\log(A\beta)$ for the AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level) is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of well-being subsistence. Binary discounting avoids such repugnancy by limiting the longevity of the dynasty. On the other hand, if the phase boundary $\log B$ for ZD is negative, a finite horizon must be optimal since an infinite horizon objective function approaches negative infinity. Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4) yield increasingly negative contributions to the population value.
With respect to the distribution of consumption among generations, the rate of compound discounting plays a key role. That is, the lower the rate of discounting is, the more equal the consumption becomes. Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in an AK setting, whereas vice versa in a ZD setting. Another concern was the optimal initial consumption in light of the well-being of future generations. In any case, the optimal initial consumption decreases in the planning horizon, and ultimately, it depends on the value of $\beta\theta$. If $\beta\theta\to 1$ (that is, with the AK/ZD setting), the optimal initial consumption approaches the level of ultimate sacrifice under an infinite planning horizon, Conversely, any finite consumption of the current generation determines the (finite) planning horizon, i.e., the size of future generations.
Appendix 1
Below, we write a Cobb–Douglas production function and its dual unit cost function:
${Y}=A{K}^{\theta}{L}^{1-\theta},$ $p={B}^{-1}{r}^{\theta}w^{1-\theta}$
where $p$, $r$, and $w$ denote prices corresponding to $Y$, $K$, and $L$, respectively. The remaining parameter $A$ is referred to as the productivity, and $B$ is the cost function-based productivity. Applying Shephard’s lemma on the dual function leads to the following.
$\frac{\partial p}{\partial r}=\frac{\theta}{B}\left(\frac{r}{w}\right)^{\theta-1}=\frac{K}{Y},$ $\frac{\partial p}{\partial w}=\frac{1-\theta}{B}\left(\frac{r}{w}\right)^{\theta}=\frac{L}{Y}$
On the basis of these equations, the marginal product of capital (MPK) can be readily evaluated as follows:
$\text{MPK}=\frac{\partial Y}{\partial K}=A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=A\theta^{\theta}(1-\theta)^{1-\theta}\left(\frac{r}{w}\right)^{\theta-1}$
$A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=\frac{Y}{K}\theta=B\left(\frac{r}{w}\right)^{\theta-1}$
where $r/w\in(0,\infty)$ is the marginal rate of substitution (MRS) between capital and labor. By comparison, we are left with $B=A\theta^{\theta}(1-\theta)^{1-\theta}$. Moreover,
$A=\frac{\text{MPK}}{\theta^{\theta}(1-\theta)^{1-\theta}}\left(\frac{r}{w}\right)^{\theta-1},$ $B=\text{MPK}\left(\frac{r}{w}\right)^{\theta-1}$
On the other hand, we recall the breakdown equation of the total output (2) and take the partial derivative as follows:
$\frac{\partial Y_{t}}{\partial K_{t}}=\left(\frac{\partial K_{t+1}}{\partial K_{t}}-1\right)+\delta=\rho+\delta=\text{MPK}$ (20)
where $\rho>0$ denotes the rate of interest. Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$.
From the study in Section 3.1, we note that the condition $A\beta<1$ implies the nonincreasing property of undiscounted contributions (or utilities less the critical level) with respect to generations. The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations. Figure 7 (left) depicts the set of possible values for $(\theta,A)$ that satisfies $A\beta<1$. Additionally, from the study in Section 3.2, we note that the condition $B<1$ implies the existence of a finite optimal planning horizon. Otherwise, if $B>1$, the optimal planning horizon will be infinite. The knife-edge case where $B=1$ implies that the optimal planning horizon is indeterminate. Figure 7 (right) depicts the set of possible values for $(\theta,A)$ that satisfies $B<1$.
Figure 7: The shaded area (marginal lines not included) corresponds to the set of parameters $(\beta,A)$ where $A\beta<1$ (left), and $(\theta,A)$ where $B=A\theta^{\theta}(1-\theta)^{1-\theta}<1$ (right).
Appendix 2
Lemma 1.
The value function of the Bellman equation (4) is as follows:
$\mathcal{V}_{t}\left[k_{t};n\right]$ $=S_{{n}-t}\theta\log(k_{t})+{R}_{t}$
where $S_{n-t}$ is specified by (6) and ${R}_{t}$ is a term that does not depend on $k_{t}$.
Proof.
We show this by induction. First, the value function at $t={n}$ is evaluated via (4):
$\mathcal{V}_{n}\left[k_{n};{n}\right]=\max_{c_{n}}\left(\log c_{n}+\beta\mathcal{V}_{n+1}[k_{{n}+1}={A}(k_{n})^{\theta}-c_{n};{n}]\right)$
Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Additionally, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, the final maximization is bounded, i.e.,
$\mathcal{V}_{n}\left[k_{n};{n}\right]=\log c_{n}=\log{A}+\theta\log k_{n}$
Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the lemma holds true for $t={n}$.
Suppose that the lemma holds true for $t+1$. Then,
$\mathcal{V}_{t}\left[k_{t};n\right]=$ $\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{V}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};n\right]\right)$
$=$ $\max_{c_{t}}\left(\log c_{t}+\beta\left(S_{{n}-t-1}\theta\log\left({A}(k_{t})^{\theta}-c_{t}\right)+{R}_{t+1}\right)\right)$
Below are the corresponding first-order condition and its solution:
$\frac{1}{c_{t}}-\frac{(\beta\theta)S_{{n}-t-1}}{{A}(k_{t})^{\theta}-c_{t}}=0,$ or, $c_{t}=\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}$ (21)
Here, we use $(\beta\theta)S_{{n}-t-1}=\beta\theta+(\beta\theta)^{2}+\cdots+(\beta\theta)^{n-t}=S_{n-t}-1$. By plugging the above solution back into the maximand, we arrive at the following result:
$\mathcal{V}_{t}[k_{t};n]$ $=\log\left(\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}\right)+\beta\left(R_{t+1}+S_{n-t-1}\theta\log\left({A}(k_{t})^{\theta}-\frac{{A}(k_{t})^{\theta}}{S_{{n}-t}}\right)\right)$
$=S_{{n}-t}\theta\log(k_{t})+\left(\beta{R}_{t+1}+S_{{n}-1}\log\left(\frac{{A}}{S_{{n}-t}}\right)+\log(S_{n-1}-1)^{S_{n-1}-1}\right)$
$=S_{{n}-t}\theta\log(k_{t})+R_{t}$
Hence, the lemma follows. ∎
Lemma 2.
The optimal trajectory of state $k_{t}^{*}[n]$ for the Bellman equation (4) is as follows:
$\log k_{t}^{*}[n]=\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}$
Proof.
We show this by induction. By plugging (21) into (2.1), we obtain:
$k_{t+1}={A}(k_{t})^{\theta}-c_{t}=\frac{S_{{n}-t}-1}{S_{{n}-t}}{A}(k_{t})^{\theta}=\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta(k_{t})^{\theta}$ (22)
As we apply $t=0$ to the above (22) and take the logarithm,
$\log k_{1}=\ln\left(\frac{S_{{n}-1}}{S_{n}}{A}\beta\theta\right)+\theta\log k_{0}$
We know that the lemma is true for $t=1$.
Suppose that the lemma is true for $t$. We then know by (22) that:
$\log k_{t+1}^{*}[n]$ $=\log\left(\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta\right)+\theta\log k_{t}^{*}[n]$
$=\log\left(\frac{S_{{n}-t-1}}{S_{{n}-t}}{A}\beta\theta\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
$=\sum_{i=1}^{t+1}\theta^{t+1-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t+1}\log k_{0}$
which indicates that the lemma is true for $t+1$. Hence, the lemma follows. ∎
Proposition 1.
The optimal consumption trajectory $c_{t}^{*}[n]$ for the Bellman equation (4) is as follows:
$\log c_{t}^{*}[n]=\log\left(\frac{{A}}{S_{{n}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{n}-i}}{S_{{n}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
Proof.
This is obvious from Lemma 2 and (21). ∎
Appendix 3
Let us evaluate the following:
$\lim_{\beta\to 1}\frac{\beta-((1-\beta){n}+1)\beta^{n+1}}{(1-\beta)^{2}}$ $=\lim_{\beta\to 1}\frac{1-(n+1)((1-\beta)n+1)\beta^{n}+n\beta^{n+1}}{-2(1-\beta)}$
$=\lim_{\beta\to 1}\frac{n(n+1)\left(2\beta^{n}-((1-\beta)n+1)\beta^{n-1}\right)}{2}$
$={n(n+1)}/{2}$
where we use L’Hôpital’s rule twice. Then, we know that
$\lim_{\beta\to 1}\frac{\beta-((1-\beta){n}+1)\beta^{n+1}}{(1-\beta)^{2}}\log(A\beta)=\frac{n(n+1)}{2}\log(A)=0$
Compliance with Ethical Standards
The authors declare that they have no conflicts of interest.