### v4

# 0000-0003-0424-585X

Abstract
[…]
journal:
1 Introduction
Standard economic models of optimal growth typically assume an infinite planning horizon. This assumption, justified by the mathematical convenience of asymptotic stability or the ethical stance of treating all future generations impartially, implicitly imposes a “perpetual existence” constraint on the dynasty (Ramsey1928). While this framework functions well in growing economies, it raises deep ethical concerns when applied to stagnant or declining economies with resource constraints. In such settings, the standard discounted utilitarian approach can lead to a “dismal” long-run equilibrium where consumption approaches zero asymptotically (DASGUPTA2019). By enforcing indefinite survival, these models may inadvertently justify the perpetuation of lives that are barely worth living, or worse, lives filled with misery—a dynamic analogue to Parfit’s “Repugnant Conclusion” in population ethics (PARFIT).
This paper revisits the problem of dynastic planning through the lens of Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005). Unlike standard utilitarianism, which sums utilities regardless of their sign, CLU introduces a normative threshold of well-being (the critical level). An individual’s existence contributes to social welfare only if their utility exceeds this critical level; otherwise, it is considered a negative contribution. By applying this ethical principle to a dynamic setting, we pose a question that is often excluded from standard growth theory: Under what economic conditions is it ethically optimal for a dynasty to continue, and when is its planned termination the morally justified conclusion of history?
Methodologically, our approach establishes a structural isomorphism between static population ethics and dynamic growth theory. In the tradition of optimal population theory initiated by Samuelson1975, a social planner typically chooses the optimal population size $N$ to maximize social welfare under resource constraints. In our dynamic framework, we reinterpret this problem by viewing the number of generations $N$ (longevity) as the choice variable. This mapping reveals that the “Repugnant Conclusion” (a large population with low quality of life) and “Infinite Misery” (an infinite dynasty with low quality of life) are two sides of the same ethical coin. Consequently, our finite-horizon solution implies a conscious decision to cease reproduction, rather than a failure of time consistency. It represents an altruistic decision to prevent the creation of future generations whose lives would fall below the critical level of well-being.
We acknowledge that other axiomatic frameworks, such as Rank-Discounted Utilitarianism (RDU), have been proposed to avoid the Repugnant Conclusion (AsheimZuber2014). RDU prevents the conclusion by strictly discounting the value of additional lives. However, this approach can lead to the opposite extreme—an “anti-population bias” or “reverse repugnant conclusion”—where a small population with high utility is preferred over a larger population that is also well-off. Furthermore, in a dynamic context, RDU implies that the welfare weight of future generations depends on their relative rank in history. This dependency destroys the recursive structure of the optimization problem and leads to time inconsistency. In contrast, CLU provides a transparent, absolute threshold for “lives worth living” that preserves both intergenerational equity (for those above the threshold) and the tractability of dynamic programming.
The remainder of the paper proceeds as follows. Section 2 describes the model of roundabout production using Cobb–Douglas technology and defines the main problem as finite horizon dynamic programming. We analytically derive the optimal consumption path and the optimal longevity $N^{*}$. In Section 3, we analyze the properties of the solution under two parameter settings: the AK setting (unity output elasticity of capital) and the Zero Discounting (ZD) setting (intergenerational equity). We show that under specific conditions—particularly when productivity is low—the ethically optimal policy is to choose a finite horizon, suggesting that sustainability does not always imply indefinite continuation. Section 4 concludes with discussions on the ethical implications of our findings for modern demographic trends.
2 Model
2.1 Critical-Level Preferences
Before describing the production technology, we first specify the social planner’s objective function. We posit a social planner whose objective is to maximize the population value based on Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005). Let $\Upsilon(c)$ denote the lifetime utility (well-being) of an individual consuming a constant level of resources $c$. The planner aims to determine the consumption path and the planning horizon (number of generations) to maximize the following social welfare function:
$\mathcal{V}=\sum_{t=0}^{N}\left(\Upsilon(c_{t})-\alpha\right)$ (1)
where $\alpha$ represents the critical level of utility. An individual’s life contributes positively to social welfare if and only if their utility exceeds this critical threshold $\alpha$; otherwise, it reduces social welfare.
For analytical tractability and to capture the essential trade-offs, we employ a logarithmic utility function of the form $\Upsilon(c)=\log c+\alpha$. This functional form is widely employed in optimal growth literature, including climate change economics (Nordhaus1992; Stern2007), which facilitates comparisons with standard benchmarks. Under this specification, the contribution of generation $t$ to the population value becomes:
$\Upsilon(c_{t})-\alpha=(\log c_{t}+\alpha)-\alpha=\log c_{t}$ (2)
Consequently, the objective function simplifies to $\mathcal{V}=\sum_{t=0}^{N}\log c_{t}$. We define the well-being subsistence level $\nu$ as the consumption level where utility is exactly zero, i.e., $\Upsilon(\nu)=0$. In our logarithmic specification, this corresponds to $\nu=e^{-\alpha}$. As illustrated in Figure 1, the domain of permissible consumption is $c>\nu$. The critical level $\alpha$ thus serves as a “buffer” above the biological subsistence level. The planner will only choose to extend the dynasty if the consumption level allows for a utility strictly greater than the critical level (i.e., $\log c_{t}>0$, or $c_{t}>1$ in normalized terms), ensuring that each generation enjoys a life worth living.
Figure 1: Left: Utility function of an individual. The consumption level for neutrality (or well-being subsistence) is denoted by $\nu$. Marginal utility and average utility coincide at $c=\omega$. Right: Contribution to the critical-level utilitarian population value, where $\alpha$ denotes the critical level. At $c=\kappa$, the utility reaches the critical level. In all cases, the domain of permissible consumption levels $c$ for existence is $c>\nu$.
2.2 Roundabout production
Let us begin by postulating a two-factor Cobb–Douglas production function characterized by constant returns to scale, as follows:
$Y_{t}={A}(K_{t})^{\theta}(L_{t})^{1-\theta}$ (3)
where $Y_{t}$ denotes the economy’s output, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all of which are effective during period $t$. For the relevant parameters, $0<\theta<1$ denotes the output elasticity of capital. The level of technology is indicated by productivity, denoted by $A$, which is assumed to be fixed throughout the considered time span. Note that ${B}\equiv{A}\theta^{\theta}(1-\theta)^{1-\theta}$ is the cost function-based productivity. [^1] Otherwise, $B$ is referred to as the dual productivity. See Appendix 1 for more details.
The breakdown of output into investment, capital depreciation, and final consumption is described below:
$Y_{t}=K_{t+1}-K_{t}+\delta K_{t}+C_{t}$ (4)
Here, $\delta\leq 1$ denotes the capital depreciation factor, and $C_{t}$ denotes aggregate consumption in period $t$. Combining equations (3) and (4) and dividing both sides by $L_{t}=L$, which we assume to be constant over time, leads to the following intertemporal dynamics (or roundaboutness) of capital intensity:
$y_{t}={A}(k_{t})^{\theta}=k_{t+1}-(1-\delta)k_{t}+c_{t}$ (5)
where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in periods $t$ and $t+1$, respectively. Additionally, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per capita output and consumption, respectively, in period $t$. The social planner aims to maximize the critical-level utilitarian population value $\mathcal{\mathcal{V}}$ as previously specified, subject to the state transition function (5), i.e.,
[…]
given the initial state $k_{0}$, we assume complete depreciation $\delta=1$ and introduce the discount factor $\beta\leq 1$. We let $\beta\to 1$ for a (critical-level) utilitarian assessment. [^2] The discount factor $\beta$ is otherwise referred to as the rate of time preference. In the context of dynastic optimization, $\beta$ may be referred to as the rate of generational preference (of the population). The optimal consumption path $(c_{0}^{*},c_{1}^{*},\cdots,c_{N}^{*})$ is clearly dependent on the planning horizon $N$. We therefore solve the above problem hierarchically. That is, we first solve for the optimal consumption path given $N$ via finite horizon dynamic programming and obtain the population value function with respect to the planning horizon $\mathcal{\mathcal{V}}[N]$; therefore, we search for the optimal planning horizon $N^{*}$.
2.3 Finite horizon dynamic programming
The primary problem here is to solve for an optimal consumption path given a planning horizon. The Bellman equation of the problem is as follows: [^3] Note that the objective function of the primary problem (6a) is given at $t=0$, i.e., $\mathcal{V}[N]=\mathcal{V}_{0}[k_{0};N]$.
$\mathcal{\mathcal{V}}_{t}\left[k_{t};\ N\right]=\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{\mathcal{V}}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};\ N\right]\right)$ (7)
The optimal trajectory of the state variable $k_{t}$ is given by Lemma 2, which we append to Appendix 2 with proofs, as follows:
$k_{t}^{*}[N]=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{N}-i}}{S_{{N}-i+1}}{A}\beta\theta\right)^{\theta^{t-i}}=(k_{0})^{\theta^{t}}\prod_{i=1}^{t}\left(\frac{S_{{N}-t+i-1}}{S_{{N}-t+i}}{A}\beta\theta\right)^{\theta^{i-1}}$ (8)
where $S_{\ell}$ is defined as follows:
$S_{\ell}\equiv\sum_{i=0}^{\ell}(\beta\theta)^{i}=1+(\beta\theta)+(\beta\theta)^{2}+\cdots+(\beta\theta)^{\ell}=\frac{1-(\beta\theta)^{\ell+1}}{1-\beta\theta}$ (9)
The following optimal trajectory of consumption is obtained by (8) and (24) Appendix 2, which must be true according to the proof of Lemma 1.
$c_{t}^{*}[N]=\frac{{A}\left(k_{t}^{*}[N]\right)^{\theta}}{S_{N-1}}=\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{N}-t}}\prod_{i=1}^{t}\left(\frac{S_{{N}-t+i-1}}{S_{{N}-t+i}}{A}\beta\theta\right)^{\theta^{i}}$ (10)
With formula (10), the population value function is given as follows:
$\mathcal{V}[N]$ $=\sum_{t=0}^{N}\beta^{t}\log c_{t}^{*}[N]$
$=\log\left(\frac{{A}(k_{0})^{\theta}}{S_{N}}\right)+\sum_{t=1}^{N}\beta^{t}\log\left(\frac{{A}(k_{0})^{\theta^{t+1}}}{S_{{N}-t}}\prod_{i=1}^{t}\left(\frac{S_{{N}-t+i-1}}{S_{{N}-t+i}}{A}\beta\theta\right)^{\theta^{i}}\right)$ (11)
We hereafter aim to maximize this function with respect to the planning horizon $N$. Our approach to analyzing the population value (11) adopts a numerical rather than an analytical framework because the derivative of $\mathcal{V}[N]$ with respect to $N$ does not seem to provide meaningful insights. In the following section, the optimal trajectory of contributions $\log c_{t}^{*}[N]$ and the population value $\mathcal{V}[N]$, for any given planning horizon $N$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$. We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon is evaluable under $\beta\theta<1$. We therefore base our study of ZD ($\beta=1$) on this parameter setting (i.e., $\beta\theta<1$, which indicates that $\theta<1$, consistent with a Cobb–Douglas model). In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$), which we term AK settings, and Cobb–Douglas production without future discounting ($\theta<1$ and $\beta=1$), which we term ZD settings. Table 1 summarizes the parameter settings chosen for the numerical examinations. Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log(A\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log B$. Cases VIII and IX correspond to the parameter settings where $\log({A}\beta)=\log B=0$.
Table 1: Parameters applied in various cases.
case A{A} $\beta$ $\theta$ $\log({A}\beta)$ $\log B$ $N^{*}$ $\mathcal{V}[N^{*}]$ $\mathcal{V}[\infty]$
I $1.012$ $0.992$ 11 ++ $\infty$ $84.7$
II $1.01$ $0.992$ 11 ++ $95$ $60.0$
III $1/\beta$ $0.992$ 11 0 $73$ $55.6$
IV $1.005$ $0.992$ 11 - $58$ $51.0$
V $1.05$ 11 $0.992$ ++ $\infty$ $\infty$
VI $1.2$ 11 *1 0 $(281)$ $41.7$
VII $1.05$ 11 $0.991$ - $117$ -$\infty$
VIII 11 11 11 0 0 $54$ $55.2$ -$\infty$
IX $1/\beta$ $0.992$ *2 0 0 $53$ $49.1$ 1843.2-1843.2
  • Note:
    For all cases I–IX, $k_{0}=150$. $B=A(1-\theta)^{1-\theta}\theta^{\theta}$ is the cost-function based productivity.
  • *1
    $\theta\approx 0.955392$ where ${1.2}(1-\theta)^{1-\theta}\theta^{\theta}=B=1$.
  • *2
    $\theta\approx 0.998982$ where $(1/0.992)(1-\theta)^{1-\theta}\theta^{\theta}=B=1$.
3 Analysis
3.1 AK setting
The AK model, which was formally developed by Frankel, is one of the simplest fundamental models of endogenous growth. Here, we employ this production model to study the population value function that discounts future generations to determine whether $N\to\infty$ is an optimal policy. The optimal consumption path (10) for the AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
$c_{t}^{*}[N]=\left(\frac{{A}k_{0}}{S_{{N}-t}}\right)\left(\frac{S_{{N}-1}}{S_{N}}{A}\beta\right)\left(\frac{S_{{N}-2}}{S_{{N}-1}}{A}\beta\right)\cdots\left(\frac{S_{{N}-t}}{S_{{N}-t+1}}{A}\beta\right)=\frac{({A}\beta)^{t}{A}k_{0}}{S_{N}}$ (12)
The population value function (11), therefore, becomes:
$\mathcal{V}[N]$ $=\sum_{t=0}^{N}\beta^{t}\log\left(\frac{({A}\beta)^{t}{A}k_{0}}{S_{N}=1+\beta+\cdots+\beta^{N}}\right)$
$=\frac{\beta-((1-\beta){N}+1)\beta^{{N}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{N}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{N}+1}}{A}k_{0}\right)$ (13)
For the sake of the analysis, let us take the derivative with respect to N{N}.
$\frac{\mathrm{d}\mathcal{V}[N]}{\mathrm{d}N}=\lambda\left(\gamma-{N}\log({A}\beta)+\log\left(1-\beta^{N+1}\right)\right)$ (14)
where $\lambda$ and $\gamma$ are given as follows:
$\lambda=\frac{\beta^{N}\log\beta^{\beta}}{1-\beta}>0,$ $\gamma=1-\frac{\log({A}\beta)^{1-\beta+\log\beta}}{\log\beta^{1-\beta}}-\log({A}(1-\beta)k_{0})$
To study the derivative sign of (14), we consider the following two functions, whose equal values, i.e., $f[N]=g[N]$, convey the first-order condition of optimality.
$f[N]=-\gamma+N\log({A}\beta),$ $g[N]=\log\left(1-\beta^{N+1}\right)$ (15)
Before we proceed, let us examine the possible range of the parameter $A$. The evaluation of the marginal product of capital (MPK) of our production function (3) net of depreciation, which should coincide with the real interest rate $\rho>0$, viz.,
$\frac{\partial Y_{t}}{\partial K_{t}}-\delta={A}\theta\left(\frac{K_{t}}{L_{t}}\right)^{\theta-1}-\delta=\rho>0$ (16)
In AK models with complete capital depreciation, $\theta=\delta=1$ leads to ${A}=1+\rho$. Therefore, assuming that A>1{A}>1 is relevant in this setting. [^4] Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross interest rate and generational preference rate of the population.
Figure 2: Left: The solid lines represent $g[N]$, as defined in (15), for cases I–IV with clear correspondences, i.e., the steepest slope corresponds to case I, while case IV corresponds to the scenario with a negative slope. The dashed line represents $f[N]$, which is also defined in (15). Right: The plots show population value functions $\mathcal{V}[N]$ for cases I–IV with obvious correspondences.
Figure 2 (left) depicts the functions $f[N]$ and $g[N]$ under different parameters for cases I–IV. Clearly, $f[N]$ is a linear function whose slope is $\log({A}\beta)$, and $g[N]$ monotonously increases and approaches zero, i.e., $g[\infty]=0$. Here, we fix the discount factor $\beta$ and differentiate the productivity A{A} at four different levels. As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[N]$ will intersect with $g[N]$ at a single point, and the population value function $\mathcal{V}[N]$ has a single peak. If ${A}>1/\beta$ so that $f[N]$ has a positive slope, then $f[N]$ and $g[N]$ could either intersect with two points where the population value $\mathcal{V}[N]$ may rise, fall and then rise again, or never intersect so that the population value may rise indefinitely with respect to the planning horizon $N$. Figure 2 (right) depicts the population value functions for cases I–IV.
To visualize the optimal trajectory of the variables in various situations, we specify them here in the form of functionals. By referencing (12), the optimal trajectory of the undiscounted contribution to the population value becomes linear with respect to $t$, as follows:
$\log c_{t}^{*}[N]=\log\left(\frac{1-\beta}{1-\beta^{N+1}}{A}k_{0}\right)+t\log({A}\beta)$ (17)
In reference to (24) Appendix 2, the optimal trajectory of capital intensity becomes as follows:
$k_{t}^{*}[N]=\frac{S_{N-t}}{{A}}c_{t}^{*}[N]=\frac{1-\beta^{N+1-t}}{1-\beta^{N+1}}({A}\beta)^{t}k_{0}$ (18)
Figure 3 displays the population value function $\mathcal{V}[N]$, specified as (13), for cases I, III, and IV on the top. [^5] These figures correspond to those depicted in Figure 2 (right). Note that, as long as $\beta<1$, the population value function will always converge to a finite value, regardless of the parameter settings in addition to $\beta$. As we let $N\to\infty$ in (13), we have the following:
$\mathcal{V}[\infty]=\frac{\log\left({A}(1-\beta)^{1-\beta}\beta^{\beta}\right)}{(1-\beta)^{2}}+\frac{\log(k_{0})}{1-\beta}$ (19)
By comparing the population values obtained at the numerical solution of the first-order condition $f[N]=g[N]$, as referenced in (15), with $\mathcal{V}[\infty]$ from (19), we know the optimal planning horizon $N^{*}$, which we display in Table 1 for cases I–IV. The middle row of Figure 3 displays the optimal trajectories of the undiscounted contributions to the population value, as described in (17), for planning horizons $N=200,400,600$ for cases I, III, and IV (from left to right). Similarly, the bottom row of Figure 3 displays the optimal trajectories of capital intensity, based on (18), for planning horizons $N=200,400,600$ for cases I, III, and IV (from left to right).
Figure 3: Population value function (top), seleted trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK setting cases I (left), III (center), and IV (right). The trajectories are selected for $N=200,400,600$.
3.2 ZD setting
Here, we study ZD, i.e., $\beta=1$, under Cobb–Douglas production with an ordinary output elasticity of capital $\theta<1$. The condition that $\beta\theta<1$ nonetheless allows us to evaluate the key summation as follows:
$S_{\infty-\tau}=\lim_{N\to\infty}\frac{1-(\beta\theta)^{N-\tau+1}}{1-\beta\theta}=\frac{1}{1-\beta\theta}$
We apply the above and $N\to\infty$ to (10) and obtain the following:
$\log c_{t}^{*}[\infty]$ $=\log\left(\frac{{A}}{S_{{\infty}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{\infty}-i}}{S_{{\infty}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
$=\log\left({{A}}{(1-\beta\theta)}\right)+\left(\theta^{t-1}+\cdots+\theta+1\right)\log({A}\beta\theta)^{\theta}+\theta^{t+1}\log k_{0}$
$=\frac{\log\left({A}(1-\beta\theta)^{1-\theta}(\beta\theta)^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\beta\theta)^{\frac{\theta}{1-\theta}}\right)$ (20)
where the optimal trajectory for undiscounted contributions is geometrically convergent. We can then apply $\beta=1$ to arrive at the following:
$\log c_{t}^{*}[\infty]=\frac{\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)}{1-\theta}+\theta^{t}\left(\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}\right)$
The corresponding population value can hence be evaluated by the infinite sum of the undiscounted contributions, i.e.,
$\mathcal{V}[\infty]=\sum_{t=0}^{\infty}\log c_{t}^{*}[\infty]=$ $\frac{\infty\log\left({A}(1-\theta)^{1-\theta}\theta^{\theta}\right)+\log(k_{0})^{\theta}-\log({A}\theta)^{\frac{\theta}{1-\theta}}}{1-\theta}$
$=$ $\begin{cases}+\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B>1\\ {\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B=1\\ -\infty&\iff{A}(1-\theta)^{1-\theta}\theta^{\theta}=B<1\end{cases}$
The above result indicates that if $B>1$, then the population value is ever increasing, and $N\to\infty$ must be the optimal solution. However, if $B<1$, then $N\to\infty$ must not be optimal, and the population value must be maximized at a finite horizon $N^{*}\ll\infty$. The case where $B=1$ is the knife-edge case in which the planning horizon $N$ does not matter (beyond a certain length) in maximizing the population value. Case V corresponds to $B>1$, and $N^{*}=\infty$; case VII corresponds to $B<1$, and $N^{*}\ll\infty$; and case VI corresponds to $B=1$, and $N^{*}$ is any number greater than 281281.666The number is subject to decimal rounding. Figure 4 depicts the population value function $\mathcal{V}[N]$, along with the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[N]$ and the capital intensity $k_{t}^{*}[N]$, given a planning horizon $N$, for sample ZD settings V, VI, and VII. The final case VIII relates to the ZD/AK setting ($\theta=\beta=1$), where $B=A=1+\rho>1$, under complete depreciation, as described in (16). Thus, if $\rho>0$, then $A>1$, in which case the population value would increase without bound, as described above, leading to $N^{*}\to\infty$. If, in turn, $A=1+\rho=1$, as in case VIII, the infinite horizon population value is evaluated as follows:
$\mathcal{V}[\infty]=\lim_{\theta\to 1}{\frac{\theta}{1-\theta}}\log\left(\frac{1-\theta}{\theta}k_{0}\right)=$-$\infty$
This indicates that there exists a finite optimum horizon $N^{*}\ll\infty$. Note that the population value function in this setting can be specified by applying a unitary discounting factor $(\beta\to 1)$ to the population value function (13) of AK models as follows:
$\mathcal{V}[N]$ $=\lim_{\beta\to 1}\left(\frac{\beta-((1-\beta){N}+1)\beta^{{N}+1}}{(1-\beta)^{2}}\log({A}\beta)+\frac{1-\beta^{{N}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{N}+1}}{A}k_{0}\right)\right)$
$=\lim_{\beta\to 1}\frac{1-\beta^{{N}+1}}{1-\beta}\log\left(\frac{1-\beta}{1-\beta^{{N}+1}}k_{0}\right)=\lim_{\beta\to 1}(N+1)\beta^{N}\log\left(\frac{k_{0}}{(N+1)\beta^{N}}\right)$
$=(N+1)\log\left(\frac{k_{0}}{N+1}\right)$
The proof for the second identity, given $A=1$, is appended to Appendix 3. The third identity is subject to L’Hôpital’s rule. By the first-order condition $\frac{\partial\mathcal{V}[N]}{\partial N}=0$, the optimal planning horizon is evaluated as $N^{*}=\exp(-1+\log k_{0})-1\approx 54.182$.
Figure 4: Population value function (top), selected trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right). The trajectories are selected for $N=200,400,600$. Note that the population value function of the top-left panel is increasing indefinitely. The parameters corresponding to each of the cases are given in Table 1.
3.3 Intergenerational inequality
Figure 3 (middle row) shows that consumption inequality across generations increases as the planning horizon extends under future discounting (except in the knife-edge case III when $\log(A\beta)=0$), whereas it seems to decrease in Figure 4 (middle row) for cases with ZD. To validate this conjecture, Figure 5 shows how the Lorenz curve shifts with respect to the planning horizon (namely, $N=200,400,600$) for AK setting cases I, III, and IV and ZD setting cases V, VI, and VII. These Lorenz curves (of Figure 5) are based on the sequence of optimal consumption levels for a given planning horizon, i.e., $c_{t}^{*}[N]$. The inequality level increases in AK settings as the planning horizon expands. On the other hand, inequality level is relatively insensitive to the planning horizon for ZD setting cases, except for case VI, when $\log B=0$, where inequality decreases as the planning horizon increases.
Figure 5: Lorenz curves for $N=200$ (solid line), $N=400$ (dotted line), and $N=600$ (dashed line). The top row panels (from left to right) correspond to AK settings I, III, and IV. The second row panels (from left to right) correspond to ZD settings V, VI, and VII.
For further analysis, let us consider the Gini index, a popular measure of inequality, defined on the basis of the optimal stream of consumption $c_{t}^{*}[N]$, as follows:
$\mathcal{G}[N]=\frac{\sum_{t^{\prime}=0}^{N}\sum_{t=0}^{N}\left|c_{t}^{*}[N]-c_{t^{\prime}}^{*}[N]\right|}{2N\sum_{t=0}^{N}c_{t}^{*}[N]}$
Figure 6 shows the Gini index $\mathcal{G}$ as a function of the planning horizon $N$ under various parameter settings. For both panels, the underlying parameters are fixed at $k_{0}=150$ and $A=1$. The left panel corresponds to the AK settings ($\theta=1$) as the discount rate is increased from $\beta=0.9$ (solid line) to $\beta=0.99$ (dashed line). In the AK settings, consumption tends to become more unequal across generations as the planning horizon expands. However, this effect is mitigated when future discounting decreases. In contrast, the right panel indicates that ZD settings tend to equalize consumption across generations as the planning horizon expands, whereas inequality is enhanced by the higher output elasticity of capital.
Figure 6: The panels depict the planning horizon vs Gini Index, i.e., $(N,\mathcal{G}[N])$ for AK setting cases with different discount rates (left), and ZD setting cases with different output elasticities of capital (right). The parameter settings for the left panel are $\theta=1$, $k_{0}=150$, $A=1$, $\beta=0.9$ (solid line), and $\beta=0.99$ (dashed line). The parameter settings for the right panel are $\beta=1$, $k_{0}=150$, $A=1$, $\theta=0.9$ (solid line), and $\theta=0.95$ (dashed line).
3.4 Initial action
Let us focus on the initial action of the optimal consumption schedule, which we specify as follows based on (10) or (24):
$c_{0}^{*}[N]=\frac{{A}(k_{0})^{\theta}}{S_{N}}=\frac{1-\beta\theta}{1-(\beta\theta)^{N}}{A}(k_{0})^{\theta}$ (21)
We take the derivative of (21) and obtain the following result:
$\frac{\mathrm{d}c_{0}^{*}[N]}{\mathrm{d}N}=\frac{A(k_{0})^{\theta}\left(1-\beta\theta\right)(\beta\theta)^{N}\log(\beta\theta)}{\left(1-(\beta\theta)^{N}\right)^{2}}<0$
since $\beta\theta<1$. In other words, the more future generations are treated equally, the more current generation must reduce their consumption. In any case, the initial action monotonically decreases in the planning horizon and ultimately converges to the following:
$c_{0}^{*}[\infty]=(1-\beta\theta){A}(k_{0})^{\theta}$ (22)
By taking its derivative with respect to $\beta$, we have:
$\frac{\partial c_{0}^{*}[\infty]}{\partial\beta}=-\theta A(k_{0})^{\theta}<0$
In other words, the more we consider future generations (by raising $\beta$), the more the current generation must reduce their consumption. From another perspective, we can solve equation $c_{0}^{*}[N]=\nu$ for $N$ with respect to (21), where $\nu$ is the well-being subsistence, as follows:
$N=\frac{\log\left((1-\beta\theta)A(k_{0})^{\theta}\right)-\log\nu}{\log(\beta\theta)}$
This $N$ is the subsistence-proof size of potential generations. Finally, for AK models, the optimal initial action by (22) is $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$, but this value approaches zero if future generations are given equal consideration as the current generation ($\beta\to 1$). That is, in AK models with an infinite planning horizon, giving ultimate consideration to future generations requires the current generation to ultimately reduce their consumption, i.e., $c_{0}^{*}[\infty]\to{\nu}$. Alternatively, if we relax the infinite horizon assumption, the initial action for AK models with ultimate consideration for the future generations becomes:
$c_{0}^{*}[N]=\lim_{\beta\to 1}\frac{1-\beta}{1-\beta^{N}}{A}k_{0}=\frac{Ak_{0}}{N}$
and we are left with the AK/ZD version of subsistence-proof size of potential generations, that is, $N=Ak_{0}/\nu$.
4 Concluding Remarks
This paper has investigated the optimal longevity of a dynasty under a Critical-Level Utilitarian framework. By treating the planning horizon as an endogenous choice variable within a dynamic programming setting, we derived closed-form analytical solutions that link economic fundamentals to the ethical duration of a society. Our analysis highlights two distinct regimes determined by productivity and time preference. In the AK setting with discounting, the sign of $\log(A\beta)$ serves as the conservative boundary determining whether the optimal horizon is finite or infinite. Similarly, in the Zero Discounting (ZD) setting, which prioritizes intergenerational equity, the sign of the cost-based productivity $\log(B)$ dictates the optimal longevity. These results demonstrate that under low productivity or strict ethical constraints, a finite horizon is not an anomaly but a structurally optimal response to avoid negative social welfare.
Our findings also reveal a striking contrast in intergenerational inequality between the two settings. In the AK setting, a longer planning horizon is associated with greater inequality across generations, although a higher discount factor $\beta$ mitigates this effect. Conversely, in the ZD setting, extending the horizon tends to reduce inequality, as the burden of capital accumulation is shared more evenly. However, a higher output elasticity of capital $\theta$ in the ZD model exacerbates long-term inequality. These observations suggest that the pursuit of longevity imposes different distributional costs depending on the underlying technology and ethical discount rates.
A key theoretical insight emerges from the behavior of the initial consumption $c_{0}$ as we approach the limit of perfect intergenerational impartiality ($\beta\to 1$). In both models, a higher $\beta$ generally requires the current generation to reduce consumption to support a longer dynasty. In the limiting case of the infinite-horizon AK model with $\beta\to 1$ (and $\theta=1$), the optimal initial consumption approaches zero (or the subsistence level). This implies that if the future holds infinite potential, the current generation is ethically compelled to make the ultimate sacrifice for the sake of an eternal dynasty. In contrast, in a stagnant economy where such sacrifice yields diminishing returns, our model prescribes termination. Thus, our framework illuminates an ethical asymmetry: it demands immense endurance when the future is promising, but offers the “mercy” of termination when the future promises only misery.
Methodologically, our approach ensures time consistency through the recursive structure of dynamic programming. Unlike ad hoc termination rules, the finite horizon derived here is the result of a consistent optimization process where the continuation value falls below the critical threshold. We acknowledge, however, that our model relies on a logarithmic utility function and a fixed population size per generation to maintain analytical tractability. Introducing more general preferences (such as CRRA utility), endogenous population size, or environmental constraints would likely preclude closed-form solutions and require numerical approaches. Exploring these complex dynamics remains a promising avenue for future research.
Finally, our results offer a normative perspective on the demographic trends observed in modern mature economies. The choice to limit the number of offspring—or in our macroeconomic interpretation, to limit the longevity of the dynasty—is often viewed as a failure of vitality. However, from the perspective of Critical-Level Utilitarianism, this may be interpreted as an altruistic termination. If the current generation anticipates that economic or environmental constraints will condemn future generations to lives barely worth living, choosing a finite horizon is a rational and ethical decision to prevent the creation of suffering.
Appendix 1
Below, we write a Cobb–Douglas production function and its dual unit cost function:
${Y}=A{K}^{\theta}{L}^{1-\theta},$ $p={B}^{-1}{r}^{\theta}w^{1-\theta}$
where $p$, $r$, and $w$ denote prices corresponding to $Y$, $K$, and $L$, respectively. The remaining parameter $A$ is referred to as the productivity, and $B$ is the cost function-based productivity. Applying Shephard’s lemma on the dual function leads to the following.
$\frac{\partial p}{\partial r}=\frac{\theta}{B}\left(\frac{r}{w}\right)^{\theta-1}=\frac{K}{Y},$ $\frac{\partial p}{\partial w}=\frac{1-\theta}{B}\left(\frac{r}{w}\right)^{\theta}=\frac{L}{Y}$
On the basis of these equations, the marginal product of capital (MPK) can be readily evaluated as follows:
$\text{MPK}=\frac{\partial Y}{\partial K}=A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=A\theta^{\theta}(1-\theta)^{1-\theta}\left(\frac{r}{w}\right)^{\theta-1}$
$A\theta\left(\frac{K}{L}\right)^{\theta-1}$ $=\frac{Y}{K}\theta=B\left(\frac{r}{w}\right)^{\theta-1}$
where $r/w\in(0,\infty)$ is the marginal rate of substitution (MRS) between capital and labor. By comparison, we are left with $B=A\theta^{\theta}(1-\theta)^{1-\theta}$. Moreover,
$A=\frac{\text{MPK}}{\theta^{\theta}(1-\theta)^{1-\theta}}\left(\frac{r}{w}\right)^{\theta-1},$ $B=\text{MPK}\left(\frac{r}{w}\right)^{\theta-1}$
On the other hand, we recall the breakdown equation of the total output (4) and take the partial derivative as follows:
$\frac{\partial Y_{t}}{\partial K_{t}}=\left(\frac{\partial K_{t+1}}{\partial K_{t}}-1\right)+\delta=\rho+\delta=\text{MPK}$ (23)
where $\rho>0$ denotes the rate of interest. Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$. From the study in Section 3.1, we note that the condition $A\beta<1$ implies the nonincreasing property of undiscounted contributions (or utilities less the critical level) with respect to generations. The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations. Figure 7 (left) depicts the set of possible values for $(\theta,A)$ that satisfies $A\beta<1$. Additionally, from the study in Section 3.2, we note that the condition $B<1$ implies the existence of a finite optimal planning horizon. Otherwise, if $B>1$, the optimal planning horizon will be infinite. The knife-edge case where $B=1$ implies that the optimal planning horizon is indeterminate. Figure 7 (right) depicts the set of possible values for $(\theta,A)$ that satisfies $B<1$.
Figure 7: The shaded area (marginal lines not included) corresponds to the set of parameters $(\beta,A)$ where $A\beta<1$ (left), and $(\theta,A)$ where $B=A\theta^{\theta}(1-\theta)^{1-\theta}<1$ (right).
Appendix 2
Lemma 1.
The value function of the Bellman equation (7) is as follows:
$\mathcal{V}_{t}\left[k_{t};N\right]$ $=S_{{N}-t}\theta\log(k_{t})+{R}_{t}$
where $S_{N-t}$ is specified by (9) and ${R}_{t}$ is a term that does not depend on $k_{t}$.
Proof.
We show this by induction. First, the value function at t=Nt={N} is evaluated via (7):
$\mathcal{V}_{N}\left[k_{N};{N}\right]=\max_{c_{N}}\left(\log c_{N}+\beta\mathcal{V}_{N+1}[k_{{N}+1}={A}(k_{N})^{\theta}-c_{N};{N}]\right)$
Since $k_{{N}+1}=0$ by (6b), $c_{N}={A}(k_{N})^{\theta}$ must be true. Additionally, $\mathcal{V}_{{N}+1}=0$ must be true for efficiency. Thus, the final maximization is bounded, i.e.,
$\mathcal{V}_{N}\left[k_{N};{N}\right]=\log c_{N}=\log{A}+\theta\log k_{N}$
Because $S_{N-N}=S_{0}=1$ and since A{A} is a constant, the lemma holds true for t=Nt={N}. Suppose that the lemma holds true for $t+1$. Then,
$\mathcal{V}_{t}\left[k_{t};N\right]=$ $\max_{c_{t}}\left(\log c_{t}+\beta\mathcal{V}_{t+1}\left[k_{t+1}={A}(k_{t})^{\theta}-c_{t};N\right]\right)$
$=$ $\max_{c_{t}}\left(\log c_{t}+\beta\left(S_{{N}-t-1}\theta\log\left({A}(k_{t})^{\theta}-c_{t}\right)+{R}_{t+1}\right)\right)$
Below are the corresponding first-order condition and its solution:
$\frac{1}{c_{t}}-\frac{(\beta\theta)S_{{N}-t-1}}{{A}(k_{t})^{\theta}-c_{t}}=0,$ or, $c_{t}=\frac{{A}(k_{t})^{\theta}}{S_{{N}-t}}$ (24)
Here, we use $(\beta\theta)S_{{N}-t-1}=\beta\theta+(\beta\theta)^{2}+\cdots+(\beta\theta)^{N-t}=S_{N-t}-1$. By plugging the above solution back into the maximand, we arrive at the following result:
$\mathcal{V}_{t}[k_{t};N]$ $=\log\left(\frac{{A}(k_{t})^{\theta}}{S_{{N}-t}}\right)+\beta\left(R_{t+1}+S_{N-t-1}\theta\log\left({A}(k_{t})^{\theta}-\frac{{A}(k_{t})^{\theta}}{S_{{N}-t}}\right)\right)$
$=S_{{N}-t}\theta\log(k_{t})+\left(\beta{R}_{t+1}+S_{{N}-1}\log\left(\frac{{A}}{S_{{N}-t}}\right)+\log(S_{N-1}-1)^{S_{N-1}-1}\right)$
$=S_{{N}-t}\theta\log(k_{t})+R_{t}$
Hence, the lemma follows. ∎
Lemma 2.
The optimal trajectory of state $k_{t}^{*}[N]$ for the Bellman equation (7) is as follows:
$\log k_{t}^{*}[N]=\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{N}-i}}{S_{{N}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}$
Proof.
We show this by induction. By plugging (24) into (6b), we obtain:
$k_{t+1}={A}(k_{t})^{\theta}-c_{t}=\frac{S_{{N}-t}-1}{S_{{N}-t}}{A}(k_{t})^{\theta}=\frac{S_{{N}-t-1}}{S_{{N}-t}}{A}\beta\theta(k_{t})^{\theta}$ (25)
As we apply $t=0$ to the above (25) and take the logarithm,
$\log k_{1}=\ln\left(\frac{S_{{N}-1}}{S_{N}}{A}\beta\theta\right)+\theta\log k_{0}$
We know that the lemma is true for $t=1$. Suppose that the lemma is true for $t$. We then know by (25) that:
$\log k_{t+1}^{*}[N]$ $=\log\left(\frac{S_{{N}-t-1}}{S_{{N}-t}}{A}\beta\theta\right)+\theta\log k_{t}^{*}[N]$
$=\log\left(\frac{S_{{N}-t-1}}{S_{{N}-t}}{A}\beta\theta\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{N}-i}}{S_{{N}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
$=\sum_{i=1}^{t+1}\theta^{t+1-i}\log\left(\frac{S_{{N}-i}}{S_{{N}-i+1}}{A}\beta\theta\right)+\theta^{t+1}\log k_{0}$
which indicates that the lemma is true for $t+1$. Hence, the lemma follows. ∎
Proposition 1.
The optimal consumption trajectory $c_{t}^{*}[N]$ for the Bellman equation (7) is as follows:
$\log c_{t}^{*}[N]=\log\left(\frac{{A}}{S_{{N}-t}}\right)+\theta\left(\sum_{i=1}^{t}\theta^{t-i}\log\left(\frac{S_{{N}-i}}{S_{{N}-i+1}}{A}\beta\theta\right)+\theta^{t}\log k_{0}\right)$
Proof.
This is obvious from Lemma 2 and (24). ∎
Appendix 3
Let us evaluate the following:
$\lim_{\beta\to 1}\frac{\beta-((1-\beta){N}+1)\beta^{N+1}}{(1-\beta)^{2}}$ $=\lim_{\beta\to 1}\frac{1-(N+1)((1-\beta)N+1)\beta^{N}+N\beta^{N+1}}{-2(1-\beta)}$
$=\lim_{\beta\to 1}\frac{N(N+1)\left(2\beta^{N}-((1-\beta)N+1)\beta^{N-1}\right)}{2}$
$={N(N+1)}/{2}$
where we use L’Hôpital’s rule twice. Then, we know that
$\lim_{\beta\to 1}\frac{\beta-((1-\beta){N}+1)\beta^{N+1}}{(1-\beta)^{2}}\log(A\beta)=\frac{N(N+1)}{2}\log(A)=0$
Compliance with Ethical Standards
The authors declare that they have no conflicts of interest.